### v1

# Qianli Wang1,2   Mingyang Wang4,5,6   Nils Feldhus1,2,8   Simon Ostermann2,3,7 Yuan Cao9   Hinrich Schütze4,6   Sebastian Möller1,2   Vera Schmitt1,2 1Quality and Usability Lab, Technische Universität Berlin  2German Research Center for Artificial Intelligence (DFKI)

Abstract
Quantization methods are widely used to accelerate inference and streamline the deployment of large language models (LLMs). While prior research has extensively investigated the degradation of various LLM capabilities due to quantization, its effects on model explainability and interpretability, which are crucial for understanding decision-making processes, remain unexplored. To address this gap, we conduct comprehensive experiments using three common quantization techniques at distinct bit widths, in conjunction with two explainability methods, counterfactual examples and natural language explanations, as well as two interpretability approaches, knowledge memorization analysis and latent multi-hop reasoning analysis. [^1] Our code will be publicly available once the paper is published. We complement our analysis with a thorough user study, evaluating selected explainability methods. Our findings reveal that, depending on the configuration, quantization can significantly impact model explainability and interpretability. Notably, the direction of this effect is not consistent, as it strongly depends on (1) the quantization method, (2) the explainability or interpretability approach, and (3) the evaluation protocol. In some settings, human evaluation shows that quantization degrades explainability, while in others, it even leads to improvements. Our work serves as a cautionary tale, demonstrating that quantization can unpredictably affect model transparency. This insight has important implications for deploying LLMs in applications where transparency is a critical requirement.
[Uncaptioned image] Through a Compressed Lens: Investigating the Impact of Quantization on LLM Explainability and Interpretability
Qianli Wang1,2   Mingyang Wang4,5,6   Nils Feldhus1,2,8   Simon Ostermann2,3,7 Yuan Cao9   Hinrich Schütze4,6   Sebastian Möller1,2   Vera Schmitt1,2 1Quality and Usability Lab, Technische Universität Berlin  2German Research Center for Artificial Intelligence (DFKI) 3Saarland Informatics Campus  4LMU Munich  5Bosch Center for Artificial Intelligence (BCAI) 6Munich Center for Machine Learning (MCML)  7Centre for European Research in Trusted AI (CERTAIN) 8BIFOLD – Berlin Institute for the Foundations of Learning and Data  9Technical University of Munich Correspondence: qianli.wang@tu-berlin.de
Figure 1: The effect of quantization on counterfactual examples (§5.1.1), natural language explanations (§5.1.1), knowledge memorization analysis (§5.1.2), and latent multi-hop reasoning (§5.1.2).
1 Introduction
With the shift towards LLMs and the focus on inference acceleration enabling more accessible and efficient deployment of LLMs, numerous quantization techniques have been developed Dettmers et al. (2022); Frantar et al. (2023); Xiao et al. (2023); Lin et al. (2024). By reducing the precision of a model’s parameters and lowering the number of bits used, quantization allows to decrease model size while mostly preserving its performance Gray and Neuhoff (1998). To date, the effects of quantization are evaluated across various aspects, including multilinguality Marchisio et al. (2024), bias Gonçalves and Strubell (2023), trustfulness Hong et al. (2024), alignment Jin et al. (2024), and instruction following capability Liu et al. (2024). Quantization has recently been ubiquitously employed for purposes of explainability and interpretability, aiming to explore black-box LLMs in areas such as parametric knowledge Namburi et al. (2023), natural language explanation Wang et al. (2024); Liu et al. (2024), and counterfactual example Bhattacharjee et al. (2024); Giorgi et al. (2025); Wang et al. (2025b). Despite the valuable insights that explainability and interpretability offer, the impact of quantization on explanation methods and explanation quality remains underexplored and has yet to be comprehensively characterized.
To bridge this gap, we present a comprehensive study that evaluates the effect of quantization on the performance and behavior of explainability and interpretability methods for LLMs. We first identify two explainability methods – counterfactual examples and natural language explanations – and two interpretability methods – knowledge memorization analysis and latent multi-hop reasoning analysis – and evaluate them on three widely used model quantization techniques (Figure 1). We conduct empirical experiments to evaluate four explainability and interpretability methods across their corresponding datasets, using four models of varying sizes in full precision and under three quantization techniques at different bit widths. We find that quantization leads to information loss within the model, and this trend becomes more pronounced in smaller models. Moreover, larger quantized models do not consistently outperform smaller full-precision models and quantization-induced degradation on explanation quality depends on the specific quantization techniques, and models deployed. Generally, integer quantization bib8 yields superior explanation quality compared to other quantization techniques across the evaluated methods. Surprisingly, lower-precision LLMs can sometimes surpass their higher-precision counterparts and a quantized model may even generate better explanations than the same model in full precision.
Secondly, we perform a user study to evaluate two selected explainability methods, assessing the explanations generated by models with different precisions for trustworthiness and coherence (Figure 1). Through the user study, we observe that explanations generated by full-precision models are perceived as slightly more trustworthy and coherent than those produced by quantized models. Meanwhile, automatic evaluation metrics for both explainability methods are not fully reliable and tend to overestimate the extent of explanation quality degradation resulting from quantization.
2 Preliminaries
Quantization.
The decoding stage during LLM inference is typically memory-bound, where the key-value cache (KV cache) overhead often exceeds the size of the model weights Li et al. (2024c). Quantization techniques compress LLMs by converting model weights, activations, or the KV cache, originally in 32-bit floating-point format, into lower-precision data types Zhu et al. (2024), e.g., 8-bit integer Dettmers et al. (2022). These techniques can be broadly categorized into two types: quantization-aware training (QAT) and post-training quantization (PTQ). QAT requires retraining to mitigate errors introduced by quantization, whereas PTQ facilitates the direct application of a quantized model during inference. Thus, PTQ is especially appealing, as it offers computational efficiency without necessitating modifications to the model architecture or training process. Among PTQ, weight-only quantization is the most conventional and widely adopted method Wan et al. (2024); Zhou et al. (2024), which effectively accelerates matrix multiplications during the decoding stage Li et al. (2024c). Thereby, in this paper, we primarily evaluate the impact of weight-only quantization (§3.3) on explainability and interpretability (§4), which eliminates the need for retraining to address errors resulting from quantization.
Impact of Quantization.
Recent work has extensively examined the impact of quantization on various capabilities of LLMs. Marchisio et al. (2024) conduct a thorough analysis of quantized multilingual LLMs, focusing on performance degradation across languages. Gonçalves and Strubell (2023); Kirsten et al. (2024) explore the emergence of bias in the outputs generated by quantized models. Liu et al. (2024) find that in-context learning ability gradually declines in heavily quantized LLMs. Jin et al. (2024) observe that models with 4-bit quantization can still retain the alignment ability. In our work, we explicitly explore the impact of quantization on both interpretability and explainability.
Explainability & Interpretability.
Both explainability and interpretability focus on understanding how LLMs arrive at their decisions and how to make those decisions more understandable to humans. Although the distinction between these terms can be subtle and they are often used interchangeably Madsen et al. (2022); Lyu et al. (2024), the nuances between them are outlined as follows: Interpretability refers to the understandability of a model’s internal structure being inherently understandable, whereas explainability involves post-hoc techniques that analyze a model’s behavior to provide insights into its decisions Zhao et al. (2024); Calderon and Reichart (2025). Hendriks et al. (2025) experiment with several distillation methods, exploring their impact on model explainability in terms of natural language explanations. Their approach to studying the effect of explainability is most aligned with our work. In this study, we investigate how various quantization methods affect model explainability and interpretability.
3 Experimental Setup
We examine four representative methods (§3.1) to evaluate the impact of quantization on explainability and interpretability. Specifically, we compare full-precision LLMs (§3.4) with LLMs quantized using different techniques and varying bit configurations (§3.3) across multiple datasets (§3.2).
3.1 Methods
3.1.1 Explainability
Counterfactual Examples (CFE).
CFEs refers to edited inputs with minimal modifications that result in different model predictions Miller (2019); Madsen et al. (2022), which can be used for improving model performance Kaushik et al. (2020) and robustness Ross et al. (2021), and understanding the black-box nature of models Wu et al. (2021). To minimize the influence of factors such as in-context learning ability degradation Liu et al. (2024), we generate CFEs in a zero-shot setting using FIZLE Bhattacharjee et al. (2024).
Natural Language Explanations (NLE).
NLEs are free-text explanations that elucidate the prediction made by the model, which humans can easily understand Camburu et al. (2018); Wiegreffe et al. (2021); Wang et al. (2025a). Analogous to the CFEs, we generate NLEs in a zero-shot setting using ZeroCoT Kojima et al. (2022).
3.1.2 Interpretability
Knowledge Memorization Analysis.
As a first interpretability technique, we investigate the memorization of a model’s factual knowledge, identifying the reasons behind potential factual forgetting Namburi et al. (2023). By leveraging the theory of knowledge neurons, which suggests that specific neurons in LLMs are responsible for storing specific pieces of knowledge, we explore how quantization alters the storage and retrieval processes within these knowledge neurons. The findings aim to shed light on the trade-offs between model efficiency and fidelity to factual information, providing deeper insights into the impact of quantization on the internal mechanisms of LLMs.
Latent Multi-hop Reasoning Analysis.
We reproduce the study on inspecting latent multi-hop reasoning errors by Yang et al. (2024). In summary, their analysis tests if LLMs use latent reasoning pathways and internal recall of knowledge about entities to answer complex factual queries. Since the authors reported a clear scaling trend for the first hop of reasoning but not for the second hop, we aim to investigate if this pattern can also be found for full-precision and quantized LLMs.
3.2 Datasets
Our study employs three widely recognized datasets [^2] Dataset examples and label distributions are detailed in Appendix A. commonly used for evaluating selected explainability and interpretability methods (§3.1).
eSNLI
Camburu et al. (2018) extends the SNLI corpus Bowman et al. (2015), which categorizes the relationship between a premise and a hypothesis into entailment, contradiction, or neutrality, by adding human-annotated NLEs.
LRE
Hernandez et al. (2024) is a knowledge probing dataset consisting of knowledge triplets $(s,r,o)$, structured in a one-hop setting, where $s$, $r$, and $o$ represents the subject, relation, and object, respectively.
TwoHop-Fact
Yang et al. (2024) is a dataset consisting of pairs of prompts: two-hop prompts ($\tau_{2H}$) for compositional queries, representing fact composition queries in the form $((e_{1},r_{1},e_{2}),(e_{2},r_{2},e_{3}))$, where $r_{1}$ and $r_{2}$ are relations and $e_{i}$ denotes an entity; and one-hop prompts ($\tau_{1H}$) for subqueries (Figure 7). $e_{1}$ and $e_{2}$ (in the second triplet) are subjects, while $e_{2}$ (in the first triplet) and $e_{3}$ objects. $e_{2}$ serves as a bridge entity, linking the two triplet to form a coherent two-hop reasoning chain.
3.3 Quantization Techniques
Building on the prior discussion (§2), we identify three commonly used PTQ techniques applied to the selected LLMs (§3.4) in our experiments:
  • GPTQ Frantar et al. (2023) uses a second-order, Hessian-based optimization to quantize weights post-training with minimal accuracy loss;
  • AWQ Lin et al. (2024) enhances weight quantization by identifying and handling activation outliers to preserve model accuracy at low bit-widths;
  • Integer quantization Dettmers et al. (2022) implemented by BitsAndBytes [^3] The implementation of integer quantization provided in BitsAndBytes is limited to weight-only quantization: https://github.com/bitsandbytes-foundation/bitsandbytes. (bib4 and bib8) enables fast and memory-efficient inference by using optimized low-bit kernels.
3.4 Models
We deploy four instruction-tuned LLMs from Qwen2.5 family Qwen et al. (2024) with model sizes ranging from 7B to 72B (Table 1), across all explainability and interpretability methods (§3.1). These models are selected because their corresponding quantized versions are officially provided [^4] https://huggingface.co/collections/Qwen/qwen25-66e81a666513e518adb90d9e, ensuring the reproducibility of our results.
4 Evaluation
4.1 Automatic Evaluation
4.1.1 Explainability
Counterfactual Examples.
We evaluate CFEs using three widely used metrics: Label flip rate (LFR), perplexity (PPL), and text similarity (TS) Nguyen et al. (2024); Wang et al. (2025b). LFR captures the frequency with which the generated CFEs alter the model prediction to a different one. PPL assesses the fluency of the generated counterfactuals by measuring how well the model predicts the next word given the previous words. As CFEs should closely resemble the original inputs, TS is calculated using normalized word-level Levenshtein distances to quantify the similarity between the generated CFEs and the original inputs.
Natural Language Explanations.
NLEs are evaluated using three metrics: Micro-$F_{1}$ Score, BARTScore, and TIGERScore. Following the predict-then-explain pattern Kumar and Talukdar (2020), Micro-$F_{1}$ Score is used to assess task performance on the target dataset. BARTScore Yuan et al. (2021) is a reference-based metric that employs BART Lewis et al. (2020) to evaluate generated explanations based on how well they align with human-annotated references. TIGERScore Jiang et al. (2024), in contrast, is a reference-free metric that deploys a fine-tuned Llama2 Touvron et al. (2023) model to identify errors in the generated explanations and assigns a penalty score between $[-5,-0.5]$ for each mistake.
4.1.2 Interpretability
Knowledge Memorization Analysis
To assess the effects of quantization on factual knowledge, we first evaluate factual recall accuracy using the LRE dataset. A comparison of results before and after quantization reveals the degree of knowledge forgetting. We then employ the knowledge attribution method (Yu and Ananiadou, 2024) to trace the information loss back to specific layers and neurons, following the knowledge neuron theory. This combined analysis uncovers how quantization impacts the internal mechanisms responsible for storing and retrieving factual information.
Latent Multi-hop Reasoning Analysis (LMHR)
[…]
4.2 Human Evaluation
We further assess the effect of quantization on explainability methods and their quality (§3.1.1) by conducting a user study in which participants subjectively evaluate the explanations (CFEs and NLEs) based on two dimensions (§4.2.1). The goal is to investigate whether the explanation quality degradation observed through automatic evaluation aligns with the findings from the user study [^5] Human evaluation of interpretability methods is omitted, as internal model states, distributed knowledge representations, and latent reasoning paths are inherently abstract. These aspects lack clear ground truth for comparison and require specialized expertise to interpret and assess the explanations being provided Alangari et al. (2023); Zhu et al. (2025)..
4.2.1 Subjective Ratings
Following the design of Likert scales for explanation evaluation proposed by Feldhus et al. (2023); Chiang and Lee (2023), and user studies on CFEs and NLEs conducted by Domnich et al. (2025); Wang et al. (2025a), we ask human annotators to evaluate CFEs and NLEs based on the following dimensions, each rated on a 5-point Likert scale ranging from "strongly disagree" (1) to "strongly agree" (5):
  • Trustworthiness: Evaluate whether the provided explanation is truthfulness and can be relied upon by humans;
  • Coherence: Assess whether the provided explanation is sensible, clear, and coherent, and effectively captures the rationale.
Model NLE CFE
Preci Micro-$F_{1}$ BART TIGER LFR $\uparrow$ PPL $\downarrow$ TS $\downarrow$
-sion Score $\uparrow$ Score $\uparrow$ Score $\uparrow$
Qwen2.5-7B full 87.40% -6.56 -0.13 82.00% 94.34 0.35
bib4 84.80% -6.69 -0.22 75.00% 64.99 2.41
bib8 86.70% -6.56 -0.24 74.20% 88.85 1.61
gptq4 86.80% -6.53 -0.51 72.00% 92.39 0.36
gptq8 87.30% -6.60 -0.16 76.60% 79.48 1.63
awq 86.10% -6.59 -0.13 74.40% 99.97 0.40
Qwen2.5-14B full 69.50% -6.52 -0.25 87.20% 90.67 0.41
bib4 63.90% -6.73 -0.24 83.20% 93.03 0.44
bib8 67.70% -6.59 -0.21 85.00% 93.54 0.41
gptq4 78.50% -6.59 -0.25 86.80% 91.45 0.41
gptq8 67.90% -6.54 -0.30 85.40% 92.58 0.41
awq 70.00% -6.57 -0.81 86.60% 91.44 0.40
Qwen2.5-32B full 88.40% -7.68 -0.44 89.00% 87.07 0.40
bib4 89.50% -10.50 -1.28 87.80% 79.35 0.47
bib8 88.10% -8.77 -0.70 86.80% 86.11 0.40
gptq4 87.90% -6.61 -0.56 89.20% 87.07 0.40
gptq8 88.60% -7.91 -0.90 87.80% 90.03 0.39
awq 90.60% -9.91 -0.27 87.20% 92.82 0.38
Qwen2.5-72B full 78.80% -6.52 -0.47 89.60% 117.85 0.31
bib4 82.00% -6.54 -0.39 89.60% 120.36 0.33
bib8 78.90% -6.51 -0.52 88.40% 112.60 0.32
gptq4 78.00% -6.52 -0.47 87.80% 119.59 0.31
gptq8 78.10% -6.55 -0.53 90.00% 115.94 0.32
awq 77.70% -6.52 -0.51 88.20% 119.61 0.32
Table 1: Automatic evaluation results of NLEs and CFEs generated using ZeroCoT and FIZLE accordingly by Qwen2.5-{7B,14B,32B,72B} with full precision and different quantization methods on eSNLI, assessed by Micro-$F_{1}$ Score, BARTScore, and TIGERScore for NLEs; label flip rate (LFR), perplexity (PPL), and text similarity (TS) for CFEs. Bold-faced values indicate the best-performing approach for each respective model, while underlined values denote the best-performing quantization method.
4.2.2 User Study Setup
We conduct a user study involving $n=48$ participants, who are all native English speakers. We randomly sample ($k=30$) dataset indices. For each subset, i.e., model-precision pair, the explanations generated by the corresponding model, either in full precision or quantized using different methods (§3.3) with Qwen2.5-{7B,32B,72B}, are evaluated by at least two human annotators. We exclude Qwen2.5-14B due to its consistently suboptimal performance, both with and without quantization. Each annotator is assigned 15 explanations, accompanied by two evaluation dimensions (§4.2.1), and tasked with assigning appropriate scores based on a given Likert scale. We report inter-annotator agreements with Krippendorff’s $\alpha$ of $0.64$ for CFEs and $0.71$ for NLEs.
5 Results
5.1 Automatic Evaluation
5.1.1 Explainability
Counterfactual Examples.
Table 1 exhibits that quantization generally affects LLMs in generating CFEs, leading to a degradation in CFEs quality, particularly in PPL, which indicates CFEs are less fluent and natural. This effect is especially pronounced for smaller LLMs, whereas larger LLMs are less impacted. Notably, in some cases, quantized models perform better or comparably to full-precision models and larger quantized LLMs overall outperform smaller full-precision LLMs.
Natural Language Explanations.
Table 1 illustrates that full-precision LLMs do not consistently outperform their corresponding quantized counterparts. LLMs with lower precision (e.g., bib4) may generate higher-quality explanations than those with higher precision (bib8). Meanwhile, degradation in TIGERScore is most noticeable in quantized models, indicating that errors occur more frequently in the NLEs generated by these models. Notably, for Qwen2.5-14B, a significant performance gap is observed in $F_{1}$ score with AWQ and in BARTScore with GPTQ. Furthermore, we find that larger quantized LLMs do not consistently outperform smaller full-precision LLMs.
Model Method Accuracy % $\uparrow$ # Correct Model Method Accuracy % $\uparrow$ # Correct
Qwen2.5-7B full 63.25 6133 Qwen2.5-32B full 72.45 7025
bib4 60.72 5887 bib4 70.47 6833
bib8 63.01 6109 bib8 72.40 7020
gptq4 60.10 5827 gptq4 71.25 6908
gptq8 63.22 6130 gptq8 72.39 7019
awq 60.60 5876 awq 71.47 6930
Qwen2.5-14B full 73.08 7086 Qwen2.5-72B full 78.26 7588
bib4 70.33 6819 bib4 76.76 7443
bib8 73.06 7084 bib8 78.22 7584
gptq4 25.20 2443 gptq4 76.77 7444
gptq8 73.03 7081 gptq8 78.32 7594
awq 70.61 6846 awq 76.72 7439
Table 2: Knowledge recall accuracy results (%) and number of correct predictions (out of 9696 queries) on the LRE dataset for Qwen2.5-{7B,14B,32B,72B} models across different quantization methods.
(a) Attention: Landmark on continent
(b) FFN: Landmark on continent
Figure 2: Layer-wise neuron contribution score drop across quantization methods for the landmark on continent relation, comparing attention sublayers (left) and feed-forward sublayers (right).
(a) Contribution score comparison
(b) Layer distribution
Figure 3: Analysis on the top-300 neurons with highest contribution scores for the landmark on continent relation under different quantization methods. (a) Average contribution scores of top 300 feed-forward neurons across different quantization methods, showing how each method affects neuron activation patterns. (b) Distribution of high-scoring neurons across model layers, showing the number of neurons that exceed the full-precision model’s 300th neuron score threshold in each layer.
(a) Qwen2.5-7B
(b) Qwen2.5-14B
Figure 4: Difference in the entity recall score (EntRec), consistency score (CnstScore) (both defined in Yang et al. (2024)), and accuracy between the full-precision and bib4 quantized models of Qwen2.5-7B and Qwen2.5-14B, evaluated across all layers.
5.1.2 Interpretability
Knowledge Memorization Analysis.
Table 2 shows that quantization introduces varying degrees of accuracy drop depending on both the model size and quantization method. Notably, the accuracy of AWQ, GPTQ4, and bib4 quantized models consistently drop across all model sizes, whereas GPTQ8 and bib8 models retain performance comparable to the original full-precision models. This effect is especially pronounced in smaller models such as Qwen2.5-{7,14}B. Furthermore, the factual recall accuracy degradation is more severe in relations where the original performance has not yet saturated, indicating that such knowledge is more fragile under quantization.
Neuron-level Attribution. Following the neuron-level method of Yu and Ananiadou (2024), we assign each neuron in Qwen2.5-7B a contribution score equal to the increase in log-probability of the correct answer token that the neuron induces. Comparing these scores before and after quantization allows us to quantify how much individual neurons’ influence changes. We analyze four relations that suffer the notable accuracy drop, landmark on continent, person father, person mother, and person sport position. Results are summarized in Figures 2 and 3; additional relations are reported in App. C.5.
Layer-wise Trends. Figure 2 depicts the drop in aggregate contribution scores for attention (left) and feed-forward (right) sub-layers. All quantization methods exhibit a pronounced decline in the final two layers, indicating that most information loss occurs during the network’s decision stages.
Model Method $r_{1}(e_{1})$uparrow $r_{2}(e_{2})$uparrow $r_{2}(r_{1}(e_{1}))$uparrow
Qwen2.5-7B full 25.03 39.07 20.61
bib4 25.01 39.02 20.61
bib8 25.01 39.02 20.61
gptq4 22.29 37.44 19.59
gptq8 24.76 39.15 20.55
awq 22.07 38.89 18.05
Qwen2.5-14B full 35.23 40.45 24.72
bib4 35.16 40.56 24.76
bib8 35.16 40.56 24.76
gptq4 32.22 45.19 23.29
gptq8 35.16 39.55 24.61
awq 24.69 35.61 21.80
Qwen2.5-32B full 21.61 36.89 18.07
bib4 21.58 36.85 18.09
bib8 21.58 36.85 18.09
gptq4 16.64 34.90 16.73
gptq8 21.18 36.59 17.70
awq 15.11 29.28 14.65
Qwen2.5-72B full 15.71 25.09 9.17
bib4 16.93 25.93 10.13
bib8 20.73 30.84 11.05
gptq4 20.11 34.96 11.14
gptq8 15.74 25.18 9.15
awq 17.08 23.65 9.10
Table 3: Accuracy (in %) of different models in connecting and traversing implicit knowledge to successfully answer latent multi-hop queries on TwoHop-Fact.
Neuron-wise Trends. We further track the 300 feed-forward neurons with the highest contribution scores in full-precision model. As shown in Figure 3, both their mean contribution and their per-layer counts decrease after quantization, with the sharpest reductions again concentrated in the last layers, mirroring the layer-level observation described above.
Collectively, these findings confirm that quantization diminishes decisive information stored in the network’s top layers, leading to the accuracy drops reported in Table 2.
Latent Multi-hop Reasoning Analysis.
Figure 4 outlines that, for Qwen2.5-{7B,14B}, quantization generally leads to weaker internal recall of the bridge entity $e_{2}$ across most layers, while the degradation of output $e_{3}$ consistency across layers remains mixed. Table 3 reveals that bib8 is the most effective quantization method at preserving performance. Meanwhile, we find that quantization substantially affects the first hop $r_{1}(e_{1})$, by as much as $30.08\%$, while its impact on the second hop is minimal, with an average of degradation of only $4.25\%$. Moreover, the $r_{2}(r_{1}(e_{1}))$ performance deterioration due to quantization is strongly correlated with the ability to correctly predict the bridge entities $r_{1}(e_{1})$, as indicated by a Spearman’s correlation of $0.93$. Exceptionally, we observe that for Qwen2.5-72B, quantization has a positive effect on overall performance, which can be attributed to the regularization effect Park et al. (2022) or the noise introduced by quantization Li et al. (2024b), as discussed in §5.1.1.
5.2 Discussion
Appendix C includes supplementary evaluations using Llama3-{8,70}B AI@Meta (2024) and LLMs (§3.4) across different datasets (§C.2) to further validate the generalizablity and robustness of our findings [^6] These experiments further evaluate CFE performance on AG News Zhang et al. (2015), focusing on news topic classification, and NLE on HealthFC Vladika et al. (2024), pertaining to fact-checking within the medical domain (App. C.2).. In general, quantization leads to information loss within the model (Figure 4), and this trend becomes more pronounced in smaller models. Perhaps surprisingly, quantization may lead to performance improvements in larger models. We observe that larger quantized models do not consistently surpass smaller full-precision models, with explanation quality contingent upon the specific quantization techniques, and deployed models. Furthermore, LLMs quantized with lower bit precision, which one would intuitively expect to exhibit reduced task performance, do not invariably perform worse than those with higher bit precision. This may be ascribable to the regularization effect Park et al. (2022) or the noise Li et al. (2024b) introduced by quantization, which limits weight precision and may inadvertently enhance the model’s robustness to slight variations. Overall, across the selected explainability and interpretability methods, bib demonstrates superiority in maximally preserving task performance and explanation quality relative to full-precision models.
5.3 User Study
Quantization NLE CFE
Model Metric Trust. Cohere. Trust. Cohere.
Qwen-7B full 3.47 3.22 3.90 3.40
gptq4 3.06 3.14 3.20 2.84
gptq8 2.96 2.63 3.26 2.86
awq 2.98 2.96 2.69 2.49
Qwen-32B full 3.41 3.41 3.50 3.21
gptq4 3.32 3.60 3.25 2.96
gptq8 3.51 3.00 3.45 3.26
awq 2.55 2.47 3.26 3.26
Qwen-72B full 3.38 3.50 4.30 4.30
gptq4 3.21 3.23 3.04 2.61
gptq8 2.97 2.86 2.92 3.02
awq 2.92 3.71 3.40 3.37
Table 4: User study results for generated NLEs and CFEs on eSNLI, evaluated based on Trustworthiness (Trust.) and Coherence (Cohere.).
Table 4 presents results from the human evaluation, showing that, overall, CFEs and NLEs generated by full-precision models are perceived as more trustworthy and coherent than those by quantized models. This can be attributed to the fact that quantization affects the model’s ability to capture contextual semantics due to distribution shifts caused by errors inevitably introduced through truncation or rounding Dettmers et al. (2022) and disrupt the model’s original confidence calibration Proskurina et al. (2024). As a result, the coherence of generated text is impaired Resendiz and Klinger (2025) and the likelihood of hallucinations increases Li et al. (2024a), ultimately diminishing annotators’ trust in the model’s explanations.
Figure 5: Explanation quality drop as measured by both automatic and human evaluation. Automatic metrics overestimate the impact of quantization on NLEs and CFEs compared to the human evaluation, averaged across three models and quantization methods (§4.2.2).
Moreover, as shown in Table 4, AWQ is less effective than GPTQ which further supports the trends observed in the automatic evaluation (§5.1.1). CFEs and NLEs produced by GPTQ4 are occasionally preferred by human annotators over those from GPTQ8, which aligns with findings from the automatic evaluation. Additionally, Figure 5 illustrates that automatic evaluations are generally unreliable in capturing the quality degradation of CFEs and NLEs due to quantization. Overall, automatic evaluation metrics for both CFEs and NLEs tend to overestimate the extent of degradation when compared to user study results. Nevertheless, we observe that automatic metrics are able to partially capture the degradation introduced by AWQ (Figure 14), showing partial alignment with the user study results.
6 Conclusion
In this work, we examine the impact of quantization on two explainability methods and two interpretability methods, evaluated on their respective datasets using three quantization techniques across four LLMs of varying sizes. We observe that, in general, quantization leads to information loss within the model, and this effect becomes more pronounced in smaller models. Moreover, larger quantized models do not necessarily outperform smaller full-precision models. The extent of explanation quality degradation is dependent on the specific quantization methods and the models employed. Lower precision LLMs can sometimes outperform their higher-precision counterparts, and quantized models may even surpass full-precision models, particularly in the case of larger LLMs. Overall, integer quantization (bib8) yields higher-quality explanations compared to other quantization techniques across the selected explainability and interpretability methods. Furthermore, a user study involving CFEs and NLEs reveals that explanations generated by quantized LLMs tend to be less coherent and trustworthy. Additionally, automatic evaluation metrics for selected explainability methods are not entirely reliable, frequently overestimating the impact of quantization on explanation quality.
Limitations
Our experimental work is confined to English-language datasets. Consequently, the effectiveness in other languages may not be comparable. Extending experiments to the multilingual setting is considered as future work.
In our experiments, we extensively compare full-precision models with different quantized versions in 4-bit and 8-bit formats. Lower-bit quantization, such as 1-bit or 2-bit, is not included in our study.
Moreover, the scope of our experiments is limited to post-training quantization techniques. Investigating the impact of weight-activation quantization, KV cache compression, or quantization-aware training techniques on explainability and interpretability is counted as future work.
As our experiments are limited to two explainability and two interpretability methods (§3.1), the findings may not fully generalize to all other explainability and interpretability approaches. We consider our work to be a first step at the emerging intersection between explainability and interpretability with model efficiency, and extending this analysis to a broader range of methods constitutes a valuable direction for future research within the community.
Ethics Statement
The participants in our user studies were compensated at or above the minimum wage in accordance with the standards of our host institutions’ regions. The annotation took each annotator 45 minutes on average.
Appendix A Dataset Information
A.1 Dataset Examples
eSNLI
Figure 6 presents an example from the eSNLI dataset consisting of premise, hypothesis, and gold label.
LRE
An exemplary query consists of $s\leftarrow\text{"{the company that created Visual Basic}"}$ and $o\leftarrow\text{"{The current CEO of}"}$. A correct answer (in this example: Satya Nadella) by the explained model is the criterion by which the LRE data is filtered.
TwoHop-Fact
Figure 7 illustrates an example from the TwoHop-Fact dataset. Based on the input text, $(e_{1},r_{1},e_{2})$ corresponds (Superstition, singer, Stevie Wonder), while $(e_{2},r_{2},e_{3})$ represents (Stevie Wonder, mother, Lula).
eSNLI (Natural Language Inference) Premise: This church choir sings to the masses as they sing joyous songs from the book at a church.
Hypothesis: The church has cracks in the ceiling.
Label: Neutral
Figure 6: An example from the eSNLI dataset.
TwoHop-Fact (Multi-hop Reasoning) One-hop prompts: ($p_{1}$): The mother of Stevie Wonder is Lula. ($p_{2}$): The singer of ‘Superstition’ is Stevie Wonder.
Two-hop prompt: ($p$): The mother of the singer of ‘Superstition’ is
Answer: Lura
Figure 7: An example from the TwoHop-Fact dataset.
A.2 Label Distribution
Label distribution of eSNLI is shown in Figure 8.
Figure 8: Label distribution of eSNLI.
Appendix B Models & Inference Time
Name Citation Size Precision Link
Llama3 AI@Meta (2024) 8B full https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct
Llama3 AI@Meta (2024) 8B GPTQ4 https://huggingface.co/TechxGenus/Meta-Llama-3-8B-Instruct-GPTQ
Llama3 AI@Meta (2024) 8B AWQ https://huggingface.co/TechxGenus/Meta-Llama-3-8B-Instruct-AWQ
Llama3 AI@Meta (2024) 70B full https://huggingface.co/meta-llama/Meta-Llama-3-70B-Instruct
Llama3 AI@Meta (2024) 70B GPTQ4 https://huggingface.co/TechxGenus/Meta-Llama-3-70B-Instruct-GPTQ
Llama3 AI@Meta (2024) 70B AWQ https://huggingface.co/TechxGenus/Meta-Llama-3-70B-Instruct-AWQ
Qwen2.5 Qwen et al. (2024) 7B Full https://huggingface.co/Qwen/Qwen2.5-7B-Instruct
Qwen2.5 Qwen et al. (2024) 7B AWQ https://huggingface.co/Qwen/Qwen2.5-7B-Instruct-AWQ
Qwen2.5 Qwen et al. (2024) 7B GPTQ4 https://huggingface.co/Qwen/Qwen2.5-7B-Instruct-GPTQ-Int4
Qwen2.5 Qwen et al. (2024) 7B GPTQ8 https://huggingface.co/Qwen/Qwen2.5-7B-Instruct-GPTQ-Int8
Qwen2.5 Qwen et al. (2024) 14B Full https://huggingface.co/Qwen/Qwen2.5-14B-Instruct
Qwen2.5 Qwen et al. (2024) 14B AWQ https://huggingface.co/Qwen/Qwen2.5-14B-Instruct-AWQ
Qwen2.5 Qwen et al. (2024) 14B GPTQ4 https://huggingface.co/Qwen/Qwen2.5-14B-Instruct-GPTQ-Int4
Qwen2.5 Qwen et al. (2024) 14B GPTQ8 https://huggingface.co/Qwen/Qwen2.5-14B-Instruct-GPTQ-Int8
Qwen2.5 Qwen et al. (2024) 32B Full https://huggingface.co/Qwen/Qwen2.5-32B-Instruct
Qwen2.5 Qwen et al. (2024) 32B AWQ https://huggingface.co/Qwen/Qwen2.5-32B-Instruct-AWQ
Qwen2.5 Qwen et al. (2024) 32B GPTQ4 https://huggingface.co/Qwen/Qwen2.5-32B-Instruct-GPTQ-Int4
Qwen2.5 Qwen et al. (2024) 32B GPTQ8 https://huggingface.co/Qwen/Qwen2.5-32B-Instruct-GPTQ-Int8
Qwen2.5 Qwen et al. (2024) 72B Full https://huggingface.co/Qwen/Qwen2.5-72B-Instruct
Qwen2.5 Qwen et al. (2024) 72B AWQ https://huggingface.co/Qwen/Qwen2.5-72B-Instruct-AWQ
Qwen2.5 Qwen et al. (2024) 72B GPTQ4 https://huggingface.co/Qwen/Qwen2.5-72B-Instruct-GPTQ-Int4
Qwen2.5 Qwen et al. (2024) 72B GPTQ8 https://huggingface.co/Qwen/Qwen2.5-72B-Instruct-GPTQ-Int8
Table 5: Detailed information about used LLMs in our experiments.
Table 5 presents details of the all LLMs used in our experiments (§3.4), including model sizes, quantization approaches and corresponding URLs from the Hugging Face Hub. All models were directly obtained from the Hugging Face repository. All experiments were conducted using A100 or H100 GPUs. Explanation generation across the entire dataset, including both natural language explanations (NLEs) and counterfactual examples (CFEs), can be completed within 10 hours.
Appendix C Additional Experiments
C.1 Models
In addition to the LLMs listed in Section 3.4, we further validate our findings using models from a different family, specifically Llama3-8B and Llama3-70B AI@Meta (2024). Detailed information about these two models is provided in Table 5.
C.2 Datasets
Additional experiments further evaluate counterfactual examples performance on AG News Zhang et al. (2015) and natural language explanations on HealthFC Vladika et al. (2024).
C.2.1 Dataset Information
AG News
Zhang et al. (2015) comprises news articles generated by merging the title and description fields from articles across four categories: World, Sports, Business, and Sci/Tech.
HealthFC
Vladika et al. (2024) is a bilingual fact-checking dataset (English and German) comprising questions, documents, veracity annotations (indicating whether the answer is true, false, or unknown based on the provided document), and corresponding human-annotated explanations.
C.2.2 Dataset Examples
Examples from the AG News and HealthFC datasets are shown in Figure 9.
AG News (News Topic Classification) News: E-mail scam targets police chief Wiltshire Police warns about ""phishing"" after its fraud squad chief was targeted.
Label: sci/tech
HealthFC (Fact-checking) Question: Does chicken soup help with colds?
Document: "However, such experiments – even though they may sound so promising – do not provide any evidence that the soup also works the same in the human body. Chicken soup in case of cold: theories without evidence Extensive studies on the effect of chicken soup do not exist, but all the more attempts to explain them. As long as there are no studies with human subjects, we simply cannot assess whether and which chicken soup ingredients could help cold-stricken people."
Label: unknown
Explanation: "So far, this has only been investigated in laboratory experiments. Studies on efficacy in humans are missing so far. Therefore, we can not judge whether chicken soup is helpful for colds."
Figure 9: Examples from AG News and HealthFC.
C.2.3 Label Distributions
Figure 10 demonstrates the label distributions of the AG News and HealthFC datasets.
(a) AG News
(b) HealthFC
Figure 10: Label distributions of AG News and HealthFC.
C.3 Natural Language Explanation
Model Dataset eSNLI HealthFC
Precision $F_{1}$ BART TIGER $F_{1}$ BART TIGER
Score $\uparrow$ Score $\uparrow$ Score $\uparrow$ Score $\uparrow$ Score $\uparrow$ Score $\uparrow$
Llama3-8B full 36.90% -6.62 -0.37 51.43% -4.29 -1.01
bib4 34.40% -6.60 -0.54 54.00% -4.65 -1.30
bib8 37.60% -6.66 -0.37 51.14% -4.26 -0.95
gptq4 38.70% -7.42 -0.29 57.71% -4.36 -1.28
awq 35.70% -6.76 -0.26 56.00% -4.43 -1.31
Llama3-70B full 57.80% -6.58 -0.16 44.00% -4.48 -1.04
bib4 59.30% -6.62 -0.71 55.71% -4.44 -1.56
bib8 43.90% -6.76 -0.26 42.86% -4.18 -1.30
gptq4 60.90% -6.62 -0.14 45.43% -4.32 -1.12
awq 61.10% -6.61 -0.15 47.71% -4.54 -1.01
Table 6: Automatic evaluation results of NLEs generated using ZeroCoT by Llama3-8B and Llama3-70B with full precision, AWQ, bib, and GPTQ quantization on eSNLI and HealthFC, assessed using Micro $F_{1}$ score, BARTScore, and TIGERScore.
Model Dataset HealthFC
Precision $F_{1}$ BART TIGER
Score $\uparrow$ Score $\uparrow$ Score $\uparrow$
Qwen2.5-7B full 39.14% -4.41 -0.34
bib4 36.57% -4.35 -0.64
bib8 36.57% -7.84 -0.94
gptq4 38.86% -4.29 -0.65
gptq8 39.71% -6.02 -0.74
awq 36.86% -4.26 -0.93
Qwen2.5-14B full 45.14% -4.25 -0.39
bib4 40.29% -4.30 -0.29
bib8 40.29% -4.34 -0.24
gptq4 36.57% -4.23 -0.41
gptq8 39.42% -4.27 -0.31
awq 46.29% -4.33 -0.48
Qwen2.5-32B full 42.57% -8.17 -2.68
bib4 42.86% -6.06 -1.61
bib8 47.43% -9.07 -3.00
gptq4 41.43% -6.62 -1.88
gptq8 42.29% -8.54 -2.86
awq 39.43% -8.97 -3.27
Qwen2.5-72B full 51.43% -4.21 -0.58
bib4 53.71% -4.21 -0.58
bib8 54.00% -4.22 -0.57
gptq4 53.42% -4.17 -0.60
gptq8 52.09% -4.20 -0.65
awq 52.28% -4.22 -0.66
Table 7: Automatic evaluation results of natural language explanations (NLEs) generated using ZeroCoT by Qwen2.5-{7,14,32,72}B with full precision and AWQ, bib, and GPTQ $\{4,8\}$-bits quantization on HealthFC, assessed by Micro $F_{1}$ score, BARTScore, and TIGERScore.
Table 6 presents automatic evaluation results of natural language explanations generated by Llama3-{8B,70B} on the eSNLI and HealthFC datasets. Table 7 provides additional evaluation results of natural language explanations generated by Vicuna-13B, and Qwen2.5-{7B,14B,32B,72B} on the HealthFC dataset. Notably, under some circumstances, quantized models outperform their full-precision counterparts. Furthermore, both tables indicate that in general, larger quantized models exhibit superior performance compared to smaller full-precision models. Additionally, integer quantization implemented by BitsAndBytes and GPTQ demonstrate compelling performance on average compared to AWQ.
C.4 Counterfactual Example
Model Dataset eSNLI AG News
Precision LFR $\uparrow$ PPL $\downarrow$ TS $\downarrow$ LFR $\uparrow$ PPL $\downarrow$ TS $\downarrow$
Llama3-8B full 90.40% 62.26 1.83 48.80% 42.88 1.54
bib4 82.00% 81.80 1.01 48.40% 46.01 1.38
bib8 86.80% 72.17 1.80 51.60% 46.95 1.60
gptq4 88.20% 76.55 1.99 65.20% 62.61 1.60
awq 87.40% 75.21 1.93 54.00% 79.87 1.21
Llama3-70B full 71.40% 80.77 0.75 49.00% 101.97 0.40
bib4 88.40% 139.80 0.58 49.60% 130.05 0.48
bib8 75.93% 102.23 1.70 26.20% 140.24 0.25
gptq4 90.40% 94.48 0.69 53.20% 115.03 0.52
awq 91.20% 114.89 0.51 47.40% 106.90 0.42
Table 8: Automatic evaluation results of CFEs generated using FIZLE by Llama3-8B and Llama3-70B with full precision, AWQ, bib, and GPTQ quantization on eSNLI and HealthFC, assessed using label flip rate (LFR), perplexity (PPL), and text similarity (TS).
Model Dataset AG News
Precision LFR $\uparrow$ PPL $\downarrow$ TS $\downarrow$
Qwen2.5-7B full 35.40% 74.87 0.53
bib4 36.00% 95.26 0.61
bib8 42.00% 72.97 0.61
gptq4 34.80% 81.52 0.57
gptq8 31.20% 79.45 0.57
awq 38.40% 80.52 0.57
Qwen2.5-14B full 38.40% 102.83 0.66
bib4 42.20% 133.47 0.78
bib8 34.60% 99.28 0.63
gptq4 39.40% 99.97 0.66
gptq8 36.80% 95.11 0.66
awq 39.20% 84.55 0.61
Qwen2.5-32B full 31.40% 83.13 0.49
bib4 39.20% 80.93 0.56
bib8 34.00% 82.82 0.49
gptq4 33.40% 81.48 0.48
gptq8 35.40% 84.81 0.49
awq 37.20% 80.60 0.52
Qwen2.5-72B full 28.80% 96.45 0.43
bib4 25.40% 93.03 0.45
bib8 31.20% 91.51 0.44
gptq4 29.60% 96.05 0.44
gptq8 26.40% 96.47 0.45
awq 24.60% 88.97 0.42
Table 9: Automatic evaluation results of CFEs generated using FIZLE by Qwen2.5-{7,14,32,72}B with full precision and AWQ, bib, and GPTQ quantization on HealthFC, assessed by label flip rate (LFR), perplexity (PPL), and text similarity (TS).
Table 8 displays the automatic evaluation results for counterfactuals generated by Llama3-{8B,70B}, assessed on the eSNLI and AG News datasets. Table 9 supplements this by presenting further evaluation results for counterfactual examples produced by Qwen2.5-{7B,14B,32B,72B}, evaluated on the AG News dataset.
Table 8 and Table 9 indicate that the quality of counterfactual examples does not uniformly enhance with larger model sizes, a finding consistent with observations shown by Nguyen et al. (2024); Bhattacharjee et al. (2024); Li et al. (2024d). We find that occasionally full-precision models outperform quantized models in generating effective counterfactual examples on AG News and eSNLI. Generally, integer quantization implemented by BitsAndBytes demonstrates superior performance compared to AWQ and GPTQ. Moreover, the conclusions of Badshah and Sajjad (2024) are not corroborated; as evidenced in Table 8 and Table 9, larger quantized models sometimes exhibit inferior performance compared to smaller full-precision models.
C.5 Knowledge Memorization Analysis
Knowledge Recall Accuracy.
Qwen2.5-7B full bib4 bib8 gptq4 gptq8 awq
city in country 93.00 85.00 93.00 93.00 93.00 93.00
company CEO 49.00 39.00 50.00 45.00 50.00 46.00
company hq 59.00 54.00 59.00 56.00 58.00 59.00
country capital city 100.00 100.00 100.00 100.00 100.00 100.00
country currency 93.00 90.00 93.00 87.00 93.00 97.00
country language 92.00 88.00 92.00 88.00 92.00 92.00
country largest city 100.00 100.00 100.00 100.00 100.00 100.00
food from country 83.00 83.00 83.00 80.00 83.00 80.00
landmark in country 81.00 80.00 81.00 79.00 81.00 80.00
landmark on continent 87.00 84.00 85.00 83.00 87.00 85.00
person father 45.00 39.00 45.00 40.00 45.00 41.00
person lead singer of band 100.00 100.00 100.00 100.00 100.00 100.00
person mother 39.00 32.00 39.00 34.00 38.00 33.00
person native language 90.00 91.00 90.00 88.00 90.00 86.00
person occupation 33.00 34.00 33.00 31.00 33.00 30.00
person plays instrument 44.00 46.00 43.00 42.00 44.00 43.00
person sport position 66.00 63.00 66.00 59.00 66.00 62.00
person university 47.00 46.00 46.00 48.00 46.00 47.00
plays pro sport 79.00 77.00 81.00 75.00 79.00 76.00
pokemon evolution 100.00 89.00 100.00 91.00 100.00 95.00
president birth year 0.00 0.00 0.00 0.00 0.00 0.00
president election year 0.00 0.00 0.00 0.00 0.00 0.00
product by company 80.00 77.00 79.00 77.00 80.00 79.00
star constellation name 85.00 85.00 85.00 85.00 85.00 84.00
superhero archnemesis 34.00 33.00 33.00 33.00 34.00 31.00
superhero person 49.00 50.00 50.00 49.00 50.00 48.00
AVG 63.25 60.72 63.01 60.10 63.22 60.00
Table 10: Per-relation knowledge recall accuracy results (%) on the LRE dataset for Qwen2.5-7B across different quantization methods.
Qwen2.5-14B full bib4 bib8 gptq4 gptq8 awq
city in country 96.00 96.00 96.00 81.00 96.00 96.00
company CEO 58.00 54.00 59.00 2.00 58.00 57.00
company hq 68.00 65.00 66.00 10.00 69.00 65.00
country capital city 100.00 100.00 100.00 79.00 100.00 100.00
country currency 90.00 83.00 87.00 77.00 87.00 90.00
country language 96.00 96.00 96.00 92.00 96.00 96.00
country largest city 96.00 100.00 96.00 83.00 96.00 96.00
food from country 80.00 77.00 80.00 50.00 80.00 80.00
landmark in country 87.00 86.00 87.00 31.00 87.00 87.00
landmark on continent 85.00 88.00 84.00 48.00 85.00 84.00
person father 63.00 56.00 62.00 7.00 63.00 57.00
person lead singer of band 100.00 100.00 100.00 29.00 100.00 100.00
person mother 59.00 49.00 59.00 1.00 59.00 53.00
person native language 94.00 94.00 94.00 39.00 93.00 93.00
person occupation 47.00 45.00 49.00 12.00 49.00 45.00
person plays instrument 67.00 61.00 68.00 52.00 66.00 63.00
person sport position 76.00 74.00 76.00 5.00 74.00 71.00
person university 48.00 45.00 52.00 43.00 47.00 46.00
plays pro sport 83.00 83.00 81.00 48.00 82.00 85.00
pokemon evolution 100.00 100.00 100.00 30.00 100.00 100.00
president birth year 0.00 0.00 0.00 0.00 0.00 0.00
president election year 0.00 0.00 0.00 0.00 0.00 0.00
product by company 85.00 82.00 85.00 38.00 85.00 84.00
star constellation name 87.00 87.00 87.00 75.00 87.00 87.00
superhero archnemesis 48.00 43.00 46.00 4.00 48.00 44.00
superhero person 76.00 73.00 75.00 5.00 77.00 72.00
AVG 73.08 70.33 73.06 25.20 73.03 70.61
Table 11: Per-relation knowledge recall accuracy results (%) on the LRE dataset for Qwen2.5-14B across different quantization methods.
Qwen2.5-32B full bib4 bib8 gptq4 gptq8 awq
city in country 85.00 85.00 85.00 85.00 85.00 85.00
company CEO 57.00 55.00 55.00 57.00 57.00 57.00
company hq 68.00 67.00 68.00 66.00 67.00 67.00
country capital city 100.00 100.00 100.00 100.00 100.00 100.00
country currency 97.00 97.00 93.00 93.00 93.00 93.00
country language 96.00 96.00 96.00 96.00 96.00 96.00
country largest city 100.00 100.00 100.00 100.00 100.00 100.00
food from country 77.00 77.00 77.00 77.00 77.00 77.00
landmark in country 87.00 86.00 86.00 87.00 87.00 87.00
landmark on continent 86.00 84.00 86.00 86.00 86.00 82.00
person father 65.00 61.00 65.00 60.00 65.00 63.00
person lead singer of band 100.00 95.00 100.00 95.00 100.00 100.00
person mother 59.00 56.00 59.00 54.00 59.00 57.00
person native language 94.00 93.00 94.00 94.00 94.00 94.00
person occupation 49.00 47.00 51.00 48.00 50.00 49.00
person plays instrument 62.00 61.00 62.00 62.00 62.00 63.00
person sport position 68.00 64.00 67.00 68.00 67.00 68.00
person university 48.00 48.00 51.00 46.00 49.00 46.00
plays pro sport 88.00 85.00 89.00 88.00 88.00 85.00
pokemon evolution 100.00 100.00 100.00 100.00 100.00 100.00
president birth year 0.00 0.00 0.00 0.00 0.00 0.00
president election year 0.00 0.00 0.00 0.00 0.00 0.00
product by company 82.00 82.00 82.00 83.00 82.00 81.00
star constellation name 89.00 87.00 89.00 88.00 89.00 89.00
superhero archnemesis 45.00 43.00 46.00 42.00 45.00 43.00
superhero person 72.00 69.00 70.00 76.00 72.00 74.00
AVG 72.45 70.47 72.40 71.25 72.39 71.47
Table 12: Per-relation knowledge recall accuracy results (%) on the LRE dataset for Qwen2.5-32B across different quantization methods.
Qwen2.5-72B full bib4 bib8 gptq4 gptq8 awq
city in country 96.00 100.00 96.00 96.00 96.00 96.00
company CEO 62.00 62.00 63.00 61.00 62.00 61.00
company hq 73.00 73.00 73.00 73.00 73.00 73.00
country capital city 100.00 100.00 100.00 100.00 100.00 100.00
country currency 80.00 80.00 80.00 80.00 80.00 83.00
country language 96.00 96.00 96.00 96.00 96.00 96.00
country largest city 100.00 100.00 100.00 100.00 100.00 100.00
food from country 83.00 83.00 83.00 83.00 83.00 83.00
landmark in country 89.00 89.00 90.00 89.00 89.00 89.00
landmark on continent 93.00 92.00 93.00 92.00 93.00 93.00
person father 74.00 71.00 74.00 69.00 74.00 71.00
person lead singer of band 90.00 100.00 90.00 95.00 90.00 100.00
person mother 69.00 65.00 69.00 64.00 70.00 66.00
person native language 93.00 93.00 93.00 92.00 93.00 93.00
person occupation 56.00 54.00 55.00 56.00 56.00 53.00
person plays instrument 73.00 69.00 73.00 71.00 73.00 69.00
person sport position 78.00 75.00 78.00 76.00 78.00 73.00
person university 52.00 49.00 53.00 49.00 52.00 52.00
plays pro sport 93.00 93.00 93.00 93.00 93.00 93.00
pokemon evolution 100.00 100.00 100.00 100.00 100.00 100.00
president birth year 0.00 0.00 0.00 0.00 0.00 0.00
president election year 0.00 0.00 0.00 0.00 0.00 0.00
product by company 85.00 85.00 85.00 85.00 85.00 84.00
star constellation name 88.00 89.00 89.00 88.00 88.00 88.00
superhero archnemesis 48.00 43.00 49.00 48.00 48.00 49.00
superhero person 78.00 80.00 78.00 79.00 78.00 77.00
AVG 78.26 76.76 78.22 76.77 78.32 76.72
Table 13: Per-relation knowledge recall accuracy results (%) on the LRE dataset for Qwen2.5-72B across different quantization methods.
In Table 10, 11, 12, 13, we provide the per-relation factual recall accuracy on Qwen2.5-{7B, 14B, 32B, 72B} models, respectively. As discussion in Section 5.1.2, we observe that the accuracy drop is more severe in relations where the original performance has not yet saturated.
Neuron Attribution Analysis.
Here we present the neuron-level knowledge attribution results on relations person father, person mother, person sport position, complementing our results in Section 5.1.2. Analysis on layer-wise neuron contribution scores are shown in Figure 11, the analysis on top-300 neurons are given in Figure 12.
(c) Attention: Person father
(d) FFN: Person father
(e) Attention: Person mother
(f) FFN: Person mother
(g) Attention: Person sport position
(h) FFN: Person sport position
Figure 11: Contribution score drop across quantization methods for other relationship types, comparing attention sublayers (left) and feed-forward sublayers (right). Rows show different relationships: (a-b) person father, (c-d) person mother, and (e-f) person sport position. The plots reveal how different quantization methods affect knowledge representation across model layers.
Refer to caption Refer to caption
(a) Person father: Contribution scores (b) Person father: Layer distribution
Refer to caption Refer to caption
(c) Person mother: Contribution scores (d) Person mother: Layer distribution
Refer to caption Refer to caption
(e) Person sport position: Contribution scores (f) Person sport position: Layer distribution
Figure 12: Analysis of other relationship types under different quantization methods. Left column: Average contribution scores of top 300 feed-forward neurons across quantization methods. Right column: Distribution of high-scoring neurons across model layers. Each row represents a different relationship: (a-b) person father, (c-d) person mother, and (e-f) person sport position. These visualizations reveal both the magnitude of contribution score changes and their distribution across the model architecture when applying different quantization techniques to various types of factual knowledge.
Appendix D Annotation
Figure 13 displays annotation guideline that we provide to human annotators. CFEs and NLEs are presented to annotators in the form of questionnaires. We use the Crowdee [^7] https://www.crowdee.com/ crowdsourcing platform to recruit annotators, distribute the questionnaires, and store their responses. A total of 48 annotators were recruited, all of whom are native English speakers without requiring specific expertise in explainable AI (XAI). Each annotators will be given 15 explanations, along with two evaluation dimensions (§4.2.1). Each explanation will be evaluated by at least two annotators.
Annotation Guideline for Human Evaluation ### User Study Description: Dear participants, Thanks for attending our user study. This study focuses on evaluating model-generated explanations. We present two types of explanations: Counterfactual Example (CFE): A minimally edited version of the input text that results in a change in the model’s prediction. Natural Language Explanation (NLE): A textual justification generated by the model to explain its decision-making process for a given input. Each explanation should be evaluated along the following two dimensions. Please assign a score to each dimension on a scale from 1 (strongly disagree) to 5 (strongly agree). Trustworthiness: Evaluate whether the provided explanation is truthfulness and can be relied upon by humans; Coherence: Assess whether the provided explanation is sensible, clear, and coherent, and effectively captures the rationale;
### Dataset Structure: e-SNLI (Stanford Natural Language Inference): Each example consists of a premise and a hypothesis. The task is to determine the relationship between the two, categorizing it as either Entailment, Contradiction, or Neutral based on the information in the premise.
e-SNLI Example: {example}
Explanation: {example explanation}
Rating:
Trustworthiness: {score}
Coherence: {score}
Entailment means the hypothesis must be true if the premise is true. Contradiction means the hypothesis must be false if the premise is true. Neutral means the hypothesis might be true, or might not — we can’t tell just from the premise.
### User Study Instruction: You will be provided with 15 instances to evaluate. For the counterfactual example evaluation, each instance includes a premise–hypothesis pair. Your task is to evaluate only the quality of the premise according to the two dimensions described above. For the natural language explanation evaluation, each instance also includes a premise–hypothesis pair, along with a model-generated justification in natural language. In this case, your task is to evaluate the provided justification based on the same two dimensions.
Figure 13: Annotation Guideline.
Appendix E Performance Degradation Comparison
Figure 14 illustrates the performance degradation between automatic and human evaluations, averaged over three automatic metrics and three human evaluation dimensions.
(a) Counterfactual Example Evaluation on Qwen2.5-7B.
(b) Counterfactual Example Evaluation on Qwen2.5-32B.
(c) Counterfactual Example Evaluation on Qwen2.5-72B.
(d) Natural Language Explanation Evaluation on Qwen2.5-7B.
(e) Natural Language Explanation Evaluation on Qwen2.5-32B.
(f) Natural Language Explanation Evaluation on Qwen2.5-72B.
Figure 14: Performance degradation between automatic and human evaluations, averaged across three automatic evaluation metrics and three human evaluation dimensions.

### v2

@@
-Qianli Wang1,2   Mingyang Wang4,5,6   Nils Feldhus1,2,8   Simon Ostermann2,3,7 Yuan Cao9   Hinrich Schütze4,6   Sebastian Möller1,2   Vera Schmitt1,2 1Quality and Usability Lab, Technische Universität Berlin  2German Research Center for Artificial Intelligence (DFKI)
+Investigating The Impact of Quantization on Factual Knowledge Recall
 Abstract
 Quantization methods are widely used to accelerate inference and streamline the deployment of large language models (LLMs).
-While prior research has extensively investigated the degradation of various LLM capabilities due to quantization, its effects on model explainability and interpretability, which are crucial for understanding decision-making processes, remain unexplored.
-To address this gap, we conduct comprehensive experiments using three common quantization techniques at distinct bit widths, in conjunction with two explainability methods, counterfactual examples and natural language explanations, as well as two interpretability approaches, knowledge memorization analysis and latent multi-hop reasoning analysis.
-[^1] Our code will be publicly available once the paper is published.
-We complement our analysis with a thorough user study, evaluating selected explainability methods.
-Our findings reveal that, depending on the configuration, quantization can significantly impact model explainability and interpretability.
-Notably, the direction of this effect is not consistent, as it strongly depends on (1) the quantization method, (2) the explainability or interpretability approach, and (3) the evaluation protocol.
-In some settings, human evaluation shows that quantization degrades explainability, while in others, it even leads to improvements.
-Our work serves as a cautionary tale, demonstrating that quantization can unpredictably affect model transparency.
-This insight has important implications for deploying LLMs in applications where transparency is a critical requirement.
-[Uncaptioned image] Through a Compressed Lens: Investigating the Impact of Quantization on LLM Explainability and Interpretability
+Although quantization’s effects on various LLM capabilities have been extensively studied, one critical area remains underexplored: factual knowledge recall (FKR), the process by which LLMs access stored knowledge.
+To this end, we conduct comprehensive experiments using three common quantization techniques at distinct bit widths, in conjunction with interpretability-driven analyses on two tasks, knowledge memorization and latent multi-hop reasoning.We show that quantization typically results in information loss within LLMs, consequently diminishing their capacity for FKR.
+This effect is particularly amplified in smaller models within the same architectural families.
+However, models quantized at reduced bit precision do not consistently exhibit inferior performance and occasionally quantization may even enhance model FKR.
+We find that BitSandBytes demonstrates highest preservation of the original full-precision model’s FKR.
+Despite variability across models and methods, quantization causes modest performance degradation and remains an effective compression strategy.
+[Uncaptioned image] Through a Compressed Lens:
+Investigating The Impact of Quantization on Factual Knowledge Recall
 Qianli Wang1,2    Mingyang Wang4,5,6    Nils Feldhus1,2,8    Simon Ostermann2,3,7 Yuan Cao9   Hinrich Schütze4,6   Sebastian Möller1,2   Vera Schmitt1,2 1Quality and Usability Lab, Technische Universität Berlin  2German Research Center for Artificial Intelligence (DFKI) 3Saarland Informatics Campus  4LMU Munich  5Bosch Center for Artificial Intelligence (BCAI) 6Munich Center for Machine Learning (MCML)  7Centre for European Research in Trusted AI (CERTAIN) 8BIFOLD – Berlin Institute for the Foundations of Learning and Data  9Technical University of Munich Correspondence: qianli.wang@tu-berlin.de
-Figure 1: The effect of quantization on counterfactual examples (§5.1.1), natural language explanations (§5.1.1), knowledge memorization analysis (§5.1.2), and latent multi-hop reasoning (§5.1.2).
 1 Introduction
-With the shift towards LLMs and the focus on inference acceleration enabling more accessible and efficient deployment of LLMs, numerous quantization techniques have been developed Dettmers et al.
-(2022); Frantar et al.
-(2023); Xiao et al.
-(2023); Lin et al.
-(2024).
-By reducing the precision of a model’s parameters and lowering the number of bits used, quantization allows to decrease model size while mostly preserving its performance Gray and Neuhoff (1998).
-To date, the effects of quantization are evaluated across various aspects, including multilinguality Marchisio et al.
-(2024), bias Gonçalves and Strubell (2023), trustfulness Hong et al.
-(2024), alignment Jin et al.
-(2024), and instruction following capability Liu et al.
-(2024).
-Quantization has recently been ubiquitously employed for purposes of explainability and interpretability, aiming to explore black-box LLMs in areas such as parametric knowledge Namburi et al.
-(2023), natural language explanation Wang et al.
-(2024); Liu et al.
-(2024), and counterfactual example Bhattacharjee et al.
-(2024); Giorgi et al.
-(2025); Wang et al.
-(2025b).
-Despite the valuable insights that explainability and interpretability offer, the impact of quantization on explanation methods and explanation quality remains underexplored and has yet to be comprehensively characterized.
-To bridge this gap, we present a comprehensive study that evaluates the effect of quantization on the performance and behavior of explainability and interpretability methods for LLMs.
-We first identify two explainability methods – counterfactual examples and natural language explanations – and two interpretability methods – knowledge memorization analysis and latent multi-hop reasoning analysis – and evaluate them on three widely used model quantization techniques (Figure 1).
-We conduct empirical experiments to evaluate four explainability and interpretability methods across their corresponding datasets, using four models of varying sizes in full precision and under three quantization techniques at different bit widths.
-We find that quantization leads to information loss within the model, and this trend becomes more pronounced in smaller models.
-Moreover, larger quantized models do not consistently outperform smaller full-precision models and quantization-induced degradation on explanation quality depends on the specific quantization techniques, and models deployed.
-Generally, integer quantization bib8 yields superior explanation quality compared to other quantization techniques across the evaluated methods.
-Surprisingly, lower-precision LLMs can sometimes surpass their higher-precision counterparts and a quantized model may even generate better explanations than the same model in full precision.
-Secondly, we perform a user study to evaluate two selected explainability methods, assessing the explanations generated by models with different precisions for trustworthiness and coherence (Figure 1).
-Through the user study, we observe that explanations generated by full-precision models are perceived as slightly more trustworthy and coherent than those produced by quantized models.
-Meanwhile, automatic evaluation metrics for both explainability methods are not fully reliable and tend to overestimate the extent of explanation quality degradation resulting from quantization.
-2 Preliminaries
+The shift towards LLMs has created strong demand for efficient inference and accessible deployment.
+In response, numerous quantization techniques have been developed dettmers-etal-2022-8bits; frantar-etal-2023-optq; xiao-etal-2023-smooth; lin-etal-2024-awq.
+By reducing the precision of a model’s parameters, quantization allows us to decrease model size while mostly preserving its performance gray-etal-1998-quantization.
+Although the effects of quantization in LLMs have been evaluated across various aspects, e.g., multilinguality, bias, fairness, and trustworthiness marchisio-etal-2024-quantization; goncalves-strubell-2023-understanding; ramesh-etal-2023-comparative; hong-etal-2024-trust; liu-etal-2024-emergent, its impact on LLMs’ ability to recall factual knowledge from pretrained memory remains underexplored and has yet to be comprehensively characterized.
+Figure 1: The effect of quantization on factual knowledge recall through knowledge memorization analysis and latent multi-hop reasoning analysis.
+To this end, we present a comprehensive study of quantization’s impact on factual knowledge recall (FKR) in LLMs.
+Instead of solely observing the performance degradation, we perform interpretability-driven analyses on knowledge memorization and latent multi-hop reasoning3.1) tasks, to scrutinize behavior across neuron, layer, and model levels and evaluate three LLMs across three widely used model quantization techniques at different bit widths (Figure 1).
+We reveal that quantization typically impairs FKR, which is particularly acute for smaller models within a given model family.
+LLMs quantized to a lower bit precision do not consistently underperform those with higher precision.
+In some cases, quantization can even paradoxically enhance factual recall.
+Our findings indicate that BitSandBytes is most effective at preserving the model’s FKR.
+We observed no significant FKR degradation from quantization that would compromise model compression effectiveness, though effects vary by model and technique.
+2 Background and Related Work
 Quantization.
-The decoding stage during LLM inference is typically memory-bound, where the key-value cache (KV cache) overhead often exceeds the size of the model weights Li et al.
-(2024c).
-Quantization techniques compress LLMs by converting model weights, activations, or the KV cache, originally in 32-bit floating-point format, into lower-precision data types Zhu et al.
-(2024), e.g., 8-bit integer Dettmers et al.
-(2022).
+Quantization techniques compress LLMs by converting model weights, activations, or the KV cache into lower-precision data types zhu-2024-surveymodelcompressionlarge.
 These techniques can be broadly categorized into two types: quantization-aware training (QAT) and post-training quantization (PTQ).
 QAT requires retraining to mitigate errors introduced by quantization, whereas PTQ facilitates the direct application of a quantized model during inference.
-Thus, PTQ is especially appealing, as it offers computational efficiency without necessitating modifications to the model architecture or training process.
-Among PTQ, weight-only quantization is the most conventional and widely adopted method Wan et al.
-(2024); Zhou et al.
-(2024), which effectively accelerates matrix multiplications during the decoding stage Li et al.
-(2024c).
-Thereby, in this paper, we primarily evaluate the impact of weight-only quantization (§3.3) on explainability and interpretability (§4), which eliminates the need for retraining to address errors resulting from quantization.
+In this paper, we primarily evaluate the impact of weight-only quantization (§3.3) on FKR (§4), which eliminates the need for retraining to address errors resulting from quantization.
 Impact of Quantization.
 Recent work has extensively examined the impact of quantization on various capabilities of LLMs.
-Marchisio et al.
-(2024) conduct a thorough analysis of quantized multilingual LLMs, focusing on performance degradation across languages.
-Gonçalves and Strubell (2023); Kirsten et al.
-(2024) explore the emergence of bias in the outputs generated by quantized models.
-Liu et al.
-(2024) find that in-context learning ability gradually declines in heavily quantized LLMs.
-Jin et al.
-(2024) observe that models with 4-bit quantization can still retain the alignment ability.
-In our work, we explicitly explore the impact of quantization on both interpretability and explainability.
-Explainability & Interpretability.
-Both explainability and interpretability focus on understanding how LLMs arrive at their decisions and how to make those decisions more understandable to humans.
-Although the distinction between these terms can be subtle and they are often used interchangeably Madsen et al.
-(2022); Lyu et al.
-(2024), the nuances between them are outlined as follows: Interpretability refers to the understandability of a model’s internal structure being inherently understandable, whereas explainability involves post-hoc techniques that analyze a model’s behavior to provide insights into its decisions Zhao et al.
-(2024); Calderon and Reichart (2025).
-Hendriks et al.
-(2025) experiment with several distillation methods, exploring their impact on model explainability in terms of natural language explanations.
-Their approach to studying the effect of explainability is most aligned with our work.
-In this study, we investigate how various quantization methods affect model explainability and interpretability.
+marchisio-etal-2024-quantization conduct a thorough analysis of quantized multilingual LLMs, focusing on performance degradation across languages.
+goncalves-strubell-2023-understanding; kirsten-etal-2024-bias explore the emergence of bias in the outputs generated by quantized models.
+liu-etal-2024-emergent find that in-context learning ability gradually declines in heavily quantized LLMs.
+jin-etal-2024-comprehensive observe that models with 4-bit quantization can still retain the alignment ability.
+singh2025interpretingeffectsquantizationllms investigate the impact of quantization on model calibration.
+In our work, we explicitly explore how quantization affects FKR.
 3 Experimental Setup
-We examine four representative methods (§3.1) to evaluate the impact of quantization on explainability and interpretability.
-Specifically, we compare full-precision LLMs (§3.4) with LLMs quantized using different techniques and varying bit configurations (§3.3) across multiple datasets (§3.2).
+We examine two representative methods (§3.1) to evaluate the impact of quantization on FKR.
+Specifically, we compare full-precision LLMs (§3.4) with LLMs quantized using different techniques and bit configurations (§3.3) across two datasets (§3.2).
 3.1 Methods
-3.1.1 Explainability
-Counterfactual Examples (CFE).
-CFEs refers to edited inputs with minimal modifications that result in different model predictions Miller (2019); Madsen et al.
-(2022), which can be used for improving model performance Kaushik et al.
-(2020) and robustness Ross et al.
-(2021), and understanding the black-box nature of models Wu et al.
-(2021).
-To minimize the influence of factors such as in-context learning ability degradation Liu et al.
-(2024), we generate CFEs in a zero-shot setting using FIZLE Bhattacharjee et al.
-(2024).
-Natural Language Explanations (NLE).
-NLEs are free-text explanations that elucidate the prediction made by the model, which humans can easily understand Camburu et al.
-(2018); Wiegreffe et al.
-(2021); Wang et al.
-(2025a).
-Analogous to the CFEs, we generate NLEs in a zero-shot setting using ZeroCoT Kojima et al.
-(2022).
-3.1.2 Interpretability
 Knowledge Memorization Analysis.
-As a first interpretability technique, we investigate the memorization of a model’s factual knowledge, identifying the reasons behind potential factual forgetting Namburi et al.
-(2023).
-By leveraging the theory of knowledge neurons, which suggests that specific neurons in LLMs are responsible for storing specific pieces of knowledge, we explore how quantization alters the storage and retrieval processes within these knowledge neurons.
-The findings aim to shed light on the trade-offs between model efficiency and fidelity to factual information, providing deeper insights into the impact of quantization on the internal mechanisms of LLMs.
+We investigate the memorization of a model’s factual knowledge, identifying the reasons behind potential factual forgetting namburi-etal-2023-cost.
+By leveraging the theory of knowledge neurons, which suggests that specific neurons in LLMs are responsible for storing specific pieces of knowledge, we explore how quantization alters the storage and retrieval processes within these neurons.
 Latent Multi-hop Reasoning Analysis.
-We reproduce the study on inspecting latent multi-hop reasoning errors by Yang et al.
-(2024).
-In summary, their analysis tests if LLMs use latent reasoning pathways and internal recall of knowledge about entities to answer complex factual queries.
-Since the authors reported a clear scaling trend for the first hop of reasoning but not for the second hop, we aim to investigate if this pattern can also be found for full-precision and quantized LLMs.
+We adapt the methodology from yang-2024-latent-multi-hop-reasoning for inspecting latent multi-hop reasoning errors.
+Specifically, the analysis tests whether full-precision and quantized LLMs employ similar latent reasoning pathways and internal knowledge recall mechanisms to answer complex factual queries.
 3.2 Datasets
-Our study employs three widely recognized datasets [^2] Dataset examples and label distributions are detailed in Appendix A.
-commonly used for evaluating selected explainability and interpretability methods (§3.1).
-eSNLI
-Camburu et al.
-(2018) extends the SNLI corpus Bowman et al.
-(2015), which categorizes the relationship between a premise and a hypothesis into entailment, contradiction, or neutrality, by adding human-annotated NLEs.
-LRE
-Hernandez et al.
-(2024) is a knowledge probing dataset consisting of knowledge triplets $(s,r,o)$, structured in a one-hop setting, where $s$, $r$, and $o$ represents the subject, relation, and object, respectively.
-TwoHop-Fact
-Yang et al.
-(2024) is a dataset consisting of pairs of prompts: two-hop prompts ($\tau_{2H}$) for compositional queries, representing fact composition queries in the form $((e_{1},r_{1},e_{2}),(e_{2},r_{2},e_{3}))$, where $r_{1}$ and $r_{2}$ are relations and $e_{i}$ denotes an entity; and one-hop prompts ($\tau_{1H}$) for subqueries (Figure 7).
+Our study employs two widely recognized datasets [^1] Dataset examples are detailed in Appendix A. for evaluating FKR with selected interpretability methods (§3.1).
+LRE hernandez-2024-linearity is a knowledge probing dataset consisting of knowledge triplets $(s,r,o)$, structured in a one-hop setting, where $s$, $r$, and $o$ represents the subject, relation, and object, respectively.
+TwoHop-Fact yang-2024-latent-multi-hop-reasoning is a dataset consisting of pairs of prompts: two-hop prompts ($\tau_{2H}$) for compositional queries, representing fact composition queries in the form $((e_{1},r_{1},e_{2}),(e_{2},r_{2},e_{3}))$, where $r_{1}$ and $r_{2}$ are relations and $e_{i}$ denotes an entity; and one-hop prompts ($\tau_{1H}$) for subqueries (Figure 4).
 $e_{1}$ and $e_{2}$ (in the second triplet) are subjects, while $e_{2}$ (in the first triplet) and $e_{3}$ objects.
@@
 Building on the prior discussion (§2), we identify three commonly used PTQ techniques applied to the selected LLMs (§3.4) in our experiments:
-  • GPTQ Frantar et al. (2023) uses a second-order, Hessian-based optimization to quantize weights post-training with minimal accuracy loss;
-  • AWQ Lin et al. (2024) enhances weight quantization by identifying and handling activation outliers to preserve model accuracy at low bit-widths;
-  • Integer quantization Dettmers et al.
-(2022) implemented by BitsAndBytes [^3] The implementation of integer quantization provided in BitsAndBytes is limited to weight-only quantization: https://github.com/bitsandbytes-foundation/bitsandbytes.
-(bib4 and bib8) enables fast and memory-efficient inference by using optimized low-bit kernels.
+  • GPTQ frantar-etal-2023-optq uses a second-order, Hessian-based optimization to quantize weights post-training with minimal accuracy loss;
+  • AWQ lin-etal-2024-awq enhances weight quantization by handling activation outliers to preserve model accuracy at low bit-widths;
+  • Integer quantization dettmers-etal-2022-8bits implemented by BitsAndBytes (bib4 and bib8) enables fast and memory-efficient inference by using optimized low-bit kernels.
 3.4 Models
-We deploy four instruction-tuned LLMs from Qwen2.5 family Qwen et al.
-(2024) with model sizes ranging from 7B to 72B (Table 1), across all explainability and interpretability methods (§3.1).
-These models are selected because their corresponding quantized versions are officially provided [^4] https://huggingface.co/collections/Qwen/qwen25-66e81a666513e518adb90d9e, ensuring the reproducibility of our results.
-4 Evaluation
-4.1 Automatic Evaluation
-4.1.1 Explainability
-Counterfactual Examples.
-We evaluate CFEs using three widely used metrics: Label flip rate (LFR), perplexity (PPL), and text similarity (TS) Nguyen et al.
-(2024); Wang et al.
-(2025b).
-LFR captures the frequency with which the generated CFEs alter the model prediction to a different one.
-PPL assesses the fluency of the generated counterfactuals by measuring how well the model predicts the next word given the previous words.
-As CFEs should closely resemble the original inputs, TS is calculated using normalized word-level Levenshtein distances to quantify the similarity between the generated CFEs and the original inputs.
-Natural Language Explanations.
-NLEs are evaluated using three metrics: Micro-$F_{1}$ Score, BARTScore, and TIGERScore.
-Following the predict-then-explain pattern Kumar and Talukdar (2020), Micro-$F_{1}$ Score is used to assess task performance on the target dataset.
-BARTScore Yuan et al.
-(2021) is a reference-based metric that employs BART Lewis et al.
-(2020) to evaluate generated explanations based on how well they align with human-annotated references.
-TIGERScore Jiang et al.
-(2024), in contrast, is a reference-free metric that deploys a fine-tuned Llama2 Touvron et al.
-(2023) model to identify errors in the generated explanations and assigns a penalty score between $[-5,-0.5]$ for each mistake.
-4.1.2 Interpretability
-Knowledge Memorization Analysis
-To assess the effects of quantization on factual knowledge, we first evaluate factual recall accuracy using the LRE dataset.
+We evaluate Llama3-8B llama3modelcard, Qwen2.5-7B and Qwen2.5-14B qwen2024qwen25technicalreport using two interpretability methods (§3.1).
+These models are selected due to the availability of quantized versions for each [^2] Detailed information for each model, including links to their respective quantized versions, is provided in Table 1., ensuring the reproducibility of our results.
+(a) Top neuron distribution (Qwen2.5-7B)
+(b) Top neuron distribution (Llama3-8B)
+(c) Attention (Qwen2.5-7B): Landmark on continent
+(d) FFN (Qwen2.5-7B): Landmark on continent
+Figure 2: Top: Distribution of high-contributing neurons across layers, showing the average number of top-300 neurons per layer for Qwen2.5-7B (left) and Llama3-8B (right).
+Bottom: Layer-wise drop in neuron contribution scores across quantization methods for the landmark on continent relation in Qwen2.5-7B, comparing attention sublayers (left) and feed-forward sublayers (right).
+4 Evaluation Setup
+Knowledge Memorization Analysis.
+To assess the effects of quantization on FKR, we first evaluate factual recall accuracy using LRE.
 A comparison of results before and after quantization reveals the degree of knowledge forgetting.
-We then employ the knowledge attribution method (Yu and Ananiadou, 2024) to trace the information loss back to specific layers and neurons, following the knowledge neuron theory.
-This combined analysis uncovers how quantization impacts the internal mechanisms responsible for storing and retrieving factual information.
-Latent Multi-hop Reasoning Analysis (LMHR)
+We then employ a knowledge attribution method (yu-ananiadou-2024-neuron) to trace the information loss back to specific layers and neurons, following the knowledge neuron theory.
+This combined analysis uncovers how quantization impacts the internal mechanisms responsible for storing and retrieving information.
+Latent Multi-hop Reasoning Analysis (LMHR).
 […]
-4.2 Human Evaluation
-We further assess the effect of quantization on explainability methods and their quality (§3.1.1) by conducting a user study in which participants subjectively evaluate the explanations (CFEs and NLEs) based on two dimensions (§4.2.1).
-The goal is to investigate whether the explanation quality degradation observed through automatic evaluation aligns with the findings from the user study [^5] Human evaluation of interpretability methods is omitted, as internal model states, distributed knowledge representations, and latent reasoning paths are inherently abstract.
-These aspects lack clear ground truth for comparison and require specialized expertise to interpret and assess the explanations being provided Alangari et al.
-(2023); Zhu et al.
-(2025)..
-4.2.1 Subjective Ratings
-Following the design of Likert scales for explanation evaluation proposed by Feldhus et al.
-(2023); Chiang and Lee (2023), and user studies on CFEs and NLEs conducted by Domnich et al.
-(2025); Wang et al.
-(2025a), we ask human annotators to evaluate CFEs and NLEs based on the following dimensions, each rated on a 5-point Likert scale ranging from "strongly disagree" (1) to "strongly agree" (5):
-  • Trustworthiness: Evaluate whether the provided explanation is truthfulness and can be relied upon by humans;
-  • Coherence: Assess whether the provided explanation is sensible, clear, and coherent, and effectively captures the rationale.
-Model NLE CFE
-Preci Micro-$F_{1}$ BART TIGER LFR $\uparrow$ PPL $\downarrow$ TS $\downarrow$
--sion Score $\uparrow$ Score $\uparrow$ Score $\uparrow$
-Qwen2.5-7B full 87.40% -6.56 -0.13 82.00% 94.34 0.35
-bib4 84.80% -6.69 -0.22 75.00% 64.99 2.41
-bib8 86.70% -6.56 -0.24 74.20% 88.85 1.61
-gptq4 86.80% -6.53 -0.51 72.00% 92.39 0.36
-gptq8 87.30% -6.60 -0.16 76.60% 79.48 1.63
-awq 86.10% -6.59 -0.13 74.40% 99.97 0.40
-Qwen2.5-14B full 69.50% -6.52 -0.25 87.20% 90.67 0.41
-bib4 63.90% -6.73 -0.24 83.20% 93.03 0.44
-bib8 67.70% -6.59 -0.21 85.00% 93.54 0.41
-gptq4 78.50% -6.59 -0.25 86.80% 91.45 0.41
-gptq8 67.90% -6.54 -0.30 85.40% 92.58 0.41
-awq 70.00% -6.57 -0.81 86.60% 91.44 0.40
-Qwen2.5-32B full 88.40% -7.68 -0.44 89.00% 87.07 0.40
-bib4 89.50% -10.50 -1.28 87.80% 79.35 0.47
-bib8 88.10% -8.77 -0.70 86.80% 86.11 0.40
-gptq4 87.90% -6.61 -0.56 89.20% 87.07 0.40
-gptq8 88.60% -7.91 -0.90 87.80% 90.03 0.39
-awq 90.60% -9.91 -0.27 87.20% 92.82 0.38
-Qwen2.5-72B full 78.80% -6.52 -0.47 89.60% 117.85 0.31
-bib4 82.00% -6.54 -0.39 89.60% 120.36 0.33
-bib8 78.90% -6.51 -0.52 88.40% 112.60 0.32
-gptq4 78.00% -6.52 -0.47 87.80% 119.59 0.31
-gptq8 78.10% -6.55 -0.53 90.00% 115.94 0.32
-awq 77.70% -6.52 -0.51 88.20% 119.61 0.32
-Table 1: Automatic evaluation results of NLEs and CFEs generated using ZeroCoT and FIZLE accordingly by Qwen2.5-{7B,14B,32B,72B} with full precision and different quantization methods on eSNLI, assessed by Micro-$F_{1}$ Score, BARTScore, and TIGERScore for NLEs; label flip rate (LFR), perplexity (PPL), and text similarity (TS) for CFEs.
-Bold-faced values indicate the best-performing approach for each respective model, while underlined values denote the best-performing quantization method.
-4.2.2 User Study Setup
-We conduct a user study involving $n=48$ participants, who are all native English speakers.
-We randomly sample ($k=30$) dataset indices.
-For each subset, i.e., model-precision pair, the explanations generated by the corresponding model, either in full precision or quantized using different methods (§3.3) with Qwen2.5-{7B,32B,72B}, are evaluated by at least two human annotators.
-We exclude Qwen2.5-14B due to its consistently suboptimal performance, both with and without quantization.
-Each annotator is assigned 15 explanations, accompanied by two evaluation dimensions (§4.2.1), and tasked with assigning appropriate scores based on a given Likert scale.
-We report inter-annotator agreements with Krippendorff’s $\alpha$ of $0.64$ for CFEs and $0.71$ for NLEs.
 5 Results
-5.1 Automatic Evaluation
-5.1.1 Explainability
-Counterfactual Examples.
-Table 1 exhibits that quantization generally affects LLMs in generating CFEs, leading to a degradation in CFEs quality, particularly in PPL, which indicates CFEs are less fluent and natural.
-This effect is especially pronounced for smaller LLMs, whereas larger LLMs are less impacted.
-Notably, in some cases, quantized models perform better or comparably to full-precision models and larger quantized LLMs overall outperform smaller full-precision LLMs.
-Natural Language Explanations.
-Table 1 illustrates that full-precision LLMs do not consistently outperform their corresponding quantized counterparts.
-LLMs with lower precision (e.g., bib4) may generate higher-quality explanations than those with higher precision (bib8).
-Meanwhile, degradation in TIGERScore is most noticeable in quantized models, indicating that errors occur more frequently in the NLEs generated by these models.
-Notably, for Qwen2.5-14B, a significant performance gap is observed in $F_{1}$ score with AWQ and in BARTScore with GPTQ.
-Furthermore, we find that larger quantized LLMs do not consistently outperform smaller full-precision LLMs.
-Model Method Accuracy % $\uparrow$ # Correct Model Method Accuracy % $\uparrow$ # Correct
-Qwen2.5-7B full 63.25 6133 Qwen2.5-32B full 72.45 7025
-bib4 60.72 5887 bib4 70.47 6833
-bib8 63.01 6109 bib8 72.40 7020
-gptq4 60.10 5827 gptq4 71.25 6908
-gptq8 63.22 6130 gptq8 72.39 7019
-awq 60.60 5876 awq 71.47 6930
-Qwen2.5-14B full 73.08 7086 Qwen2.5-72B full 78.26 7588
-bib4 70.33 6819 bib4 76.76 7443
-bib8 73.06 7084 bib8 78.22 7584
-gptq4 25.20 2443 gptq4 76.77 7444
-gptq8 73.03 7081 gptq8 78.32 7594
-awq 70.61 6846 awq 76.72 7439
-Table 2: Knowledge recall accuracy results (%) and number of correct predictions (out of 9696 queries) on the LRE dataset for Qwen2.5-{7B,14B,32B,72B} models across different quantization methods.
-(a) Attention: Landmark on continent
-(b) FFN: Landmark on continent
-Figure 2: Layer-wise neuron contribution score drop across quantization methods for the landmark on continent relation, comparing attention sublayers (left) and feed-forward sublayers (right).
+5.1 Knowledge Memorization Analysis
+Table 2 shows that quantization introduces varying degrees of accuracy drop depending on both the model size and quantization method.
+Notably, the accuracy drop is more obvious in smaller models within the same model family.
+Additionally, the accuracy of AWQ, GPTQ4, and bib4 quantized models consistently drop across all model sizes, whereas GPTQ8 and bib8 models retain performance comparable to the full-precision models.
+Furthermore, relations where full-precision performance has not saturated tend to exhibit more severe factual knowledge recall degradation (Table 4), indicating that such knowledge is more fragile under quantization.
+Neuron-level Trends.
+Following the neuron-level method of yu-ananiadou-2024-neuron, we assign each neuron in the model a contribution score equal to the increase in log-probability of the correct answer token that the neuron induces.
+Comparing these scores before and after quantization allows us to quantify how much individual neurons’ influence changes.
+We analyze four relations that suffer a notable accuracy drop: landmark on continent, person father, person mother, and person sport position.
+We track the top-300 feed-forward neurons in full-precision model, set $\tau$ as their minimum contribution score, and count how many neurons in quantized models exceed $\tau$.
+As shown in Figure 2 and 2, the per-layer counts decrease after quantization for both models, with the most obvious reductions in the last layers.
+Complete results across relations are provided in Figures 7 and 8 in App.
+C.1.
+Layer-wise Trends.
+Figure 2 depicts the drop in aggregate contribution scores for attention (left) and feed-forward (right) sub-layers.
+All quantization methods exhibit a pronounced decline in the final two layers on Qwen2.5-7B, this trend is consistent on all relations we investigate, as shown in Figure 5 in App C.1.
+While on Llama3-8B (see Figure 6), the decline occurs in middle-to-late layers, with an increase in the final layers, indicating that information loss patterns vary across model architectures.
+This divergence may be ascribable to the different ways in which factual knowledge is stored across model families choe2025autoregressivetransformersrememberfacts.
+Collectively, both layer-level and neuron-level analyses reveal that quantization primarily affects the network’s last layers.
+These findings confirm that quantization degrades the decisive information stored in these late layers, accounting for the factual knowledge recall degradation reported in Table 2.333The complete analysis of both models across different relations is provided in Appendix C.1.
+(a) Qwen2.5-7B
+(b) Qwen2.5-14B
+(c) Llama3-8B
+Figure 3: Difference in the entity recall score (EntRec), consistency score (CnstScore), and accuracy between the GPTQ8 quantized and full-precision models, evaluated across all layers.
+5.2 Latent Multi-hop Reasoning Analysis
+Quantization affects the first-hop Reasoning the most.
+Table 3 reveals that quantization substantially affects the first hop $r_{1}(e_{1})$, by as much as $30.08\%$, while its impact on the second hop is minimal, with an average of degradation of only $4.25\%$.
+The $r_{2}(r_{1}(e_{1}))$ deterioration due to quantization is strongly correlated with the ability to correctly predict the bridge entities $r_{1}(e_{1})$, as indicated by a Spearman’s correlation of $0.93$.
+Nevertheless, the FKR deterioration is not dramatic and considerably acceptable (e.g., Qwen2.5-7B shows a minor average deterioration of 0.77%).
+Quantization Effects are not consistent.
+Figure 3 illustrates that quantization effects on FKR are largely unpredictable.
+The quantization effect is heterogeneous across layers and variable across models, particularly across architectures, given that facts are stored in various ways across Llama3 and Qwen2.5 models choe2025autoregressivetransformersrememberfacts.
+Besides, different quantization approaches often affect the model FKR in different manners (Figure 12).
+Nevertheless, for a given quantization method with different bit width, the layer-wise impact remains broadly similar.
+Surprisingly, quantization can occasionally even improve FKR (Figure 3).
+This effect may be attributed to the regularization effect park-etal-2022-robust or quantization-induced noise li2024investigating, which can inadvertently enhance the model’s capability to recall factual knowledge.
+bib largely preserve the factual knowledge recall capability.
+For Qwen2.5-7B, quantization consistently reduce knowledge recall, with the largest degradations under GPTQ and AWQ (Figure 12).
+By contrast, bib4/8 degrades one-hop reasoning $r_{1}(e_{1})$ and $r_{2}(e_{2})$, as reflected by modest shifts in EntRec and CnstScore relative to the full-precision model, but leaves two-hop reasoning largely unaffected.
+For Qwen2.5-14B, bib4/8 likewise outperforms other quantization methods, best preserving and occasionally improving FKR (Figure 14).
+Moreover, despite its greater layer depth than Qwen2.5-7B, Qwen2.5-14B exhibits a similar layer-wise pattern of bib4/8 quantization effects (Figure 12(e), 14(e)), whereas other quantization methods do not.
+For Llama3-8B, the effect of different quantization methods are alike, i.e., EntRec and CnstScore of quantized model are lower than full-precision model in shallow layers, but becomes much higher in deeper layers (Figure 15).
+Among the three LLMs evaluated, Llama3-8B is least affected by quantization in terms of FKR.
+6 Conclusion
+In this work, we examined the impact of quantization on factual knowledge recall.
+In general, quantization leads to information loss within the model, which in turn degrades factual knowledge recall; this trend becomes more pronounced for smaller models within each model family.
+Moreover, LLMs quantized with lower bit precision do not invariably perform worse than those with higher bit precision.
+Quantization can occasionally even improve factual knowledge recall.
+While quantization effects vary by model and technique, we observed no performance degradation severe enough to compromise its viability as a compression strategy.
+Limitations
+Our experimental work is confined to English-language datasets.
+Consequently, the effectiveness of our experiments in other languages may not be comparable and multilingual factual knowledge recall may be simultaneously affected by the degradation of multilingual capabilities due to quantization marchisio-etal-2024-quantization.
+Extending experiments to the multilingual settings is considered for future work.
+We restrict our experiments to Llama3-8B, Qwen2.5-7B and Qwen2.5-14B for computational feasibility: our gradient-based analyses exceed the available GPU memory for larger models.
+Between the model families (Qwen, Llama, Mistral, DeepSeek), there are lots of architectural equivalences and similarities, e.g., the same attention (grouped-query attention), position embeddings (RoPE), normalization (RMSNorm) or FFN activation (SwiGLU).
+We argue that, based on our comprehensive experiments, our results are considerably robust and generalizable given the similar architectures compared to other model families.
+In our experiments, we extensively compare full-precision models with different quantized versions in 4-bit and 8-bit formats.
+Lower-bit quantization, such as 1-bit or 2-bit, is not included in our study.
+Moreover, the scope of our experiments is limited to post-training quantization techniques.
+Investigating the impact of weight-activation quantization, KV cache compression, or quantization-aware training techniques on factual knowledge recall is counted as future work.
+Appendix A Dataset Information
+LRE
+An exemplary query consists of $s\leftarrow\text{"{the company that created Visual Basic}"}$ and $o\leftarrow\text{"{The current CEO of}"}$.
+A correct answer (in this example: Satya Nadella) by the explained model is the criterion by which the LRE data is filtered.
+TwoHop-Fact
+Figure 4 illustrates an example from the TwoHop-Fact dataset.
+Based on the input text, $(e_{1},r_{1},e_{2})$ corresponds (Superstition, singer, Stevie Wonder), while $(e_{2},r_{2},e_{3})$ represents (Stevie Wonder, mother, Lula).
+TwoHop-Fact (Multi-hop Reasoning) One-hop prompts: ($p_{1}$): The mother of Stevie Wonder is Lula.
+($p_{2}$): The singer of ‘Superstition’ is Stevie Wonder.
+Two-hop prompt: ($p$): The mother of the singer of ‘Superstition’ is Answer: Lura
+Figure 4: An example from the TwoHop-Fact dataset.
+Appendix B Models & Inference Time
+Name Citation Size Precision Link
+Llama3 llama3modelcard 8B full https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct
+Llama3 llama3modelcard 8B GPTQ4 https://huggingface.co/TechxGenus/Meta-Llama-3-8B-Instruct-GPTQ
+Llama3 llama3modelcard 8B AWQ https://huggingface.co/TechxGenus/Meta-Llama-3-8B-Instruct-AWQ
+Qwen2.5 qwen2024qwen25technicalreport 7B Full https://huggingface.co/Qwen/Qwen2.5-7B-Instruct
+Qwen2.5 qwen2024qwen25technicalreport 7B AWQ https://huggingface.co/Qwen/Qwen2.5-7B-Instruct-AWQ
+Qwen2.5 qwen2024qwen25technicalreport 7B GPTQ4 https://huggingface.co/Qwen/Qwen2.5-7B-Instruct-GPTQ-Int4
+Qwen2.5 qwen2024qwen25technicalreport 7B GPTQ8 https://huggingface.co/Qwen/Qwen2.5-7B-Instruct-GPTQ-Int8
+Qwen2.5 qwen2024qwen25technicalreport 14B Full https://huggingface.co/Qwen/Qwen2.5-14B-Instruct
+Qwen2.5 qwen2024qwen25technicalreport 14B AWQ https://huggingface.co/Qwen/Qwen2.5-14B-Instruct-AWQ
+Qwen2.5 qwen2024qwen25technicalreport 14B GPTQ4 https://huggingface.co/Qwen/Qwen2.5-14B-Instruct-GPTQ-Int4
+Qwen2.5 qwen2024qwen25technicalreport 14B GPTQ8 https://huggingface.co/Qwen/Qwen2.5-14B-Instruct-GPTQ-Int8
+Table 1: Detailed information about used LLMs in our experiments.
+Table 1 presents details of the all LLMs used in our experiments (§3.4), including model sizes, quantization approaches and corresponding URLs from the Hugging Face Hub.
+All models were directly obtained from the Hugging Face repository.
+All experiments were conducted using A100 or H100 GPUs.
+Neuron-level and layer-level attribution can be completed within 10 hours, while LMHR experiments take 30 hours averagely.
+Appendix C Additional Experiments
+C.1 Knowledge Memorization Analysis
+Model Method Accuracy % $\uparrow$ # Correct
+Qwen2.5-7B full 63.25 6133
+bib4 60.72 5887
+bib8 63.01 6109
+gptq4 60.10 5827
+gptq8 63.22 6130
+awq 60.60 5876
+Qwen2.5-14B full 73.08 7086
+bib4 70.33 6819
+bib8 73.06 7084
+gptq4 25.20 2443
+gptq8 73.03 7081
+awq 70.61 6846
+Llama3-8B full 77.62 7526
+bib4 72.19 7000
+bib8 76.95 7461
+gptq8 71.39 6922
+awq 71.83 6965
+Table 2: Knowledge recall accuracy results (%) and number of correct predictions (out of 9696 queries) on the LRE dataset for Qwen2.5-{7B,14B} and Llama3-8B models across different quantization methods.
+Table 2 illustrates knowledge recall accuracy results on the LRE dataset for Qwen2.5-{7B,14B} and Llama3-8B models across different quantization methods.
+Knowledge Recall Accuracy.
+In Table 4, and 5, we provide the per-relation factual recall accuracy on Qwen2.5-7B, Qwen2.5-14B, and Llama3-8B models, respectively.
+As discussion in Section 5.1, we observe that the accuracy drop is more severe in relations where the original performance has not yet saturated.
+Neuron Attribution Analysis.
+Here we present the neuron-level knowledge attribution results on relations person father, person mother, person sport position, complementing our results in Section 5.1.
+Analysis on layer-wise neuron contribution scores are shown in Figure 5 and Figure 6, the analysis on top-300 neurons are given in Figure 9 and Figure 10.
+(a) Attention: Person father
+(b) FFN: Person father
+(c) Attention: Person mother
+(d) FFN: Person mother
+(e) Attention: Person sport position
+(f) FFN: Person sport position
+Figure 5: Contribution score drop across quantization methods on Qwen2.5-7B for other relationship types, comparing attention sublayers (left) and feed-forward sublayers (right).
+Rows show different relationships: (a-b) person father, (c-d) person mother, and (e-f) person sport position.
+The plots reveal how different quantization methods affect knowledge representation across model layers.
+(a) Attention: Person father
+(b) FFN: Person father
+(c) Attention: Person mother
+(d) FFN: Person mother
+(e) Attention: Person sport position
+(f) FFN: Person sport position
+Figure 6: Contribution score drop across quantization methods on Llama3-8B for other relationship types, comparing attention sublayers (left) and feed-forward sublayers (right).
+Rows show different relationships: (a-b) person father, (c-d) person mother, and (e-f) person sport position.
+The plots reveal how different quantization methods affect knowledge representation across model layers.
 (a) Contribution score comparison
 (b) Layer distribution
-Figure 3: Analysis on the top-300 neurons with highest contribution scores for the landmark on continent relation under different quantization methods.
+Figure 7: Analysis of the top-300 neurons with highest contribution scores for the landmark on continent relation under different quantization methods applied to Qwen2.5-7B.
 (a) Average contribution scores of top 300 feed-forward neurons across different quantization methods, showing how each method affects neuron activation patterns.
-(b) Distribution of high-scoring neurons across model layers, showing the number of neurons that exceed the full-precision model’s 300th neuron score threshold in each layer.
-(a) Qwen2.5-7B
-(b) Qwen2.5-14B
-Figure 4: Difference in the entity recall score (EntRec), consistency score (CnstScore) (both defined in Yang et al.
-(2024)), and accuracy between the full-precision and bib4 quantized models of Qwen2.5-7B and Qwen2.5-14B, evaluated across all layers.
-5.1.2 Interpretability
-Knowledge Memorization Analysis.
-Table 2 shows that quantization introduces varying degrees of accuracy drop depending on both the model size and quantization method.
-Notably, the accuracy of AWQ, GPTQ4, and bib4 quantized models consistently drop across all model sizes, whereas GPTQ8 and bib8 models retain performance comparable to the original full-precision models.
-This effect is especially pronounced in smaller models such as Qwen2.5-{7,14}B.
-Furthermore, the factual recall accuracy degradation is more severe in relations where the original performance has not yet saturated, indicating that such knowledge is more fragile under quantization.
-Neuron-level Attribution.
-Following the neuron-level method of Yu and Ananiadou (2024), we assign each neuron in Qwen2.5-7B a contribution score equal to the increase in log-probability of the correct answer token that the neuron induces.
-Comparing these scores before and after quantization allows us to quantify how much individual neurons’ influence changes.
-We analyze four relations that suffer the notable accuracy drop, landmark on continent, person father, person mother, and person sport position.
-Results are summarized in Figures 2 and 3; additional relations are reported in App.
-C.5.
-Layer-wise Trends.
-Figure 2 depicts the drop in aggregate contribution scores for attention (left) and feed-forward (right) sub-layers.
-All quantization methods exhibit a pronounced decline in the final two layers, indicating that most information loss occurs during the network’s decision stages.
-Model Method $r_{1}(e_{1})$uparrow $r_{2}(e_{2})$uparrow $r_{2}(r_{1}(e_{1}))$uparrow
+(b) Distribution of high-scoring neurons across layers, showing the number of neurons exceeding the full-precision model’s 300th neuron score threshold in each layer.
+(a) Contribution score comparison
+(b) Layer distribution
+Figure 8: Analysis of the top-300 neurons with highest contribution scores for the landmark on continent relation under different quantization methods applied to Llama3-8B.
+(a) Average contribution scores of top 300 feed-forward neurons across different quantization methods, showing how each method affects neuron activation patterns.
+(b) Distribution of high-scoring neurons across layers, showing the number of neurons exceeding the full-precision model’s 300th neuron score threshold in each layer.
+Refer to caption Refer to caption
+(a) Person father: Contribution scores (b) Person father: Layer distribution
+Refer to caption Refer to caption
+(c) Person mother: Contribution scores (d) Person mother: Layer distribution
+Refer to caption Refer to caption
+(e) Person sport position: Contribution scores (f) Person sport position: Layer distribution
+Figure 9: Analysis of other relationship types under different quantization methods applied to Qwen2.5-7B.
+Left column: Average contribution scores of top 300 feed-forward neurons across quantization methods.
+Right column: Distribution of high-scoring neurons across model layers.
+Each row represents a different relationship: (a-b) person father, (c-d) person mother, and (e-f) person sport position.
+These visualizations reveal both the magnitude of contribution score changes and their distribution across the model architecture when applying different quantization techniques to various types of factual knowledge.
+Refer to caption Refer to caption
+(a) Person father: Contribution scores (b) Person father: Layer distribution
+Refer to caption Refer to caption
+(c) Person mother: Contribution scores (d) Person mother: Layer distribution
+Refer to caption Refer to caption
+(e) Person sport position: Contribution scores (f) Person sport position: Layer distribution
+Figure 10: Analysis of other relationship types under different quantization methods applied to Llama3-8B.
+Left column: Average contribution scores of top 300 feed-forward neurons across quantization methods.
+Right column: Distribution of high-scoring neurons across model layers.
+Each row represents a different relationship: (a-b) person father, (c-d) person mother, and (e-f) person sport position.
+These visualizations reveal both the magnitude of contribution score changes and their distribution across the model architecture when applying different quantization techniques to various types of factual knowledge.
+C.2 Latent Multi-hop Reasoning
+Model Method $r_{1}(e_{1})$ $\uparrow$ $r_{2}(e_{2})$ $\uparrow$ $r_{2}(r_{1}(e_{1}))$ $\uparrow$
 Qwen2.5-7B full 25.03 39.07 20.61
@@
 awq 24.69 35.61 21.80
-Qwen2.5-32B full 21.61 36.89 18.07
-bib4 21.58 36.85 18.09
-bib8 21.58 36.85 18.09
-gptq4 16.64 34.90 16.73
-gptq8 21.18 36.59 17.70
-awq 15.11 29.28 14.65
-Qwen2.5-72B full 15.71 25.09 9.17
-bib4 16.93 25.93 10.13
-bib8 20.73 30.84 11.05
-gptq4 20.11 34.96 11.14
-gptq8 15.74 25.18 9.15
-awq 17.08 23.65 9.10
+Llama3-8B full 7.79 21.39 4.45
+bib4 4.94 27.44 3.14
+bib8 5.79 18.27 2.90
+gptq8 23.62 40.73 20.94
+awq 22.35 37.79 19.56
 Table 3: Accuracy (in %) of different models in connecting and traversing implicit knowledge to successfully answer latent multi-hop queries on TwoHop-Fact.
-Neuron-wise Trends.
-We further track the 300 feed-forward neurons with the highest contribution scores in full-precision model.
-As shown in Figure 3, both their mean contribution and their per-layer counts decrease after quantization, with the sharpest reductions again concentrated in the last layers, mirroring the layer-level observation described above.
-Collectively, these findings confirm that quantization diminishes decisive information stored in the network’s top layers, leading to the accuracy drops reported in Table 2.
-Latent Multi-hop Reasoning Analysis.
-Figure 4 outlines that, for Qwen2.5-{7B,14B}, quantization generally leads to weaker internal recall of the bridge entity $e_{2}$ across most layers, while the degradation of output $e_{3}$ consistency across layers remains mixed.
-Table 3 reveals that bib8 is the most effective quantization method at preserving performance.
-Meanwhile, we find that quantization substantially affects the first hop $r_{1}(e_{1})$, by as much as $30.08\%$, while its impact on the second hop is minimal, with an average of degradation of only $4.25\%$.
-Moreover, the $r_{2}(r_{1}(e_{1}))$ performance deterioration due to quantization is strongly correlated with the ability to correctly predict the bridge entities $r_{1}(e_{1})$, as indicated by a Spearman’s correlation of $0.93$.
-Exceptionally, we observe that for Qwen2.5-72B, quantization has a positive effect on overall performance, which can be attributed to the regularization effect Park et al.
-(2022) or the noise introduced by quantization Li et al.
-(2024b), as discussed in §5.1.1.
-5.2 Discussion
-Appendix C includes supplementary evaluations using Llama3-{8,70}B AI@Meta (2024) and LLMs (§3.4) across different datasets (§C.2) to further validate the generalizablity and robustness of our findings [^6] These experiments further evaluate CFE performance on AG News Zhang et al.
-(2015), focusing on news topic classification, and NLE on HealthFC Vladika et al.
-(2024), pertaining to fact-checking within the medical domain (App.
-C.2)..
-In general, quantization leads to information loss within the model (Figure 4), and this trend becomes more pronounced in smaller models.
-Perhaps surprisingly, quantization may lead to performance improvements in larger models.
-We observe that larger quantized models do not consistently surpass smaller full-precision models, with explanation quality contingent upon the specific quantization techniques, and deployed models.
-Furthermore, LLMs quantized with lower bit precision, which one would intuitively expect to exhibit reduced task performance, do not invariably perform worse than those with higher bit precision.
-This may be ascribable to the regularization effect Park et al.
-(2022) or the noise Li et al.
-(2024b) introduced by quantization, which limits weight precision and may inadvertently enhance the model’s robustness to slight variations.
-Overall, across the selected explainability and interpretability methods, bib demonstrates superiority in maximally preserving task performance and explanation quality relative to full-precision models.
-5.3 User Study
-Quantization NLE CFE
-Model Metric Trust. Cohere. Trust. Cohere.
-Qwen-7B full 3.47 3.22 3.90 3.40
-gptq4 3.06 3.14 3.20 2.84
-gptq8 2.96 2.63 3.26 2.86
-awq 2.98 2.96 2.69 2.49
-Qwen-32B full 3.41 3.41 3.50 3.21
-gptq4 3.32 3.60 3.25 2.96
-gptq8 3.51 3.00 3.45 3.26
-awq 2.55 2.47 3.26 3.26
-Qwen-72B full 3.38 3.50 4.30 4.30
-gptq4 3.21 3.23 3.04 2.61
-gptq8 2.97 2.86 2.92 3.02
-awq 2.92 3.71 3.40 3.37
-Table 4: User study results for generated NLEs and CFEs on eSNLI, evaluated based on Trustworthiness (Trust.) and Coherence (Cohere.).
-Table 4 presents results from the human evaluation, showing that, overall, CFEs and NLEs generated by full-precision models are perceived as more trustworthy and coherent than those by quantized models.
-This can be attributed to the fact that quantization affects the model’s ability to capture contextual semantics due to distribution shifts caused by errors inevitably introduced through truncation or rounding Dettmers et al.
-(2022) and disrupt the model’s original confidence calibration Proskurina et al.
-(2024).
-As a result, the coherence of generated text is impaired Resendiz and Klinger (2025) and the likelihood of hallucinations increases Li et al.
-(2024a), ultimately diminishing annotators’ trust in the model’s explanations.
-Figure 5: Explanation quality drop as measured by both automatic and human evaluation.
-Automatic metrics overestimate the impact of quantization on NLEs and CFEs compared to the human evaluation, averaged across three models and quantization methods (§4.2.2).
-Moreover, as shown in Table 4, AWQ is less effective than GPTQ which further supports the trends observed in the automatic evaluation (§5.1.1).
-CFEs and NLEs produced by GPTQ4 are occasionally preferred by human annotators over those from GPTQ8, which aligns with findings from the automatic evaluation.
-Additionally, Figure 5 illustrates that automatic evaluations are generally unreliable in capturing the quality degradation of CFEs and NLEs due to quantization.
-Overall, automatic evaluation metrics for both CFEs and NLEs tend to overestimate the extent of degradation when compared to user study results.
-Nevertheless, we observe that automatic metrics are able to partially capture the degradation introduced by AWQ (Figure 14), showing partial alignment with the user study results.
-6 Conclusion
-In this work, we examine the impact of quantization on two explainability methods and two interpretability methods, evaluated on their respective datasets using three quantization techniques across four LLMs of varying sizes.
-We observe that, in general, quantization leads to information loss within the model, and this effect becomes more pronounced in smaller models.
-Moreover, larger quantized models do not necessarily outperform smaller full-precision models.
-The extent of explanation quality degradation is dependent on the specific quantization methods and the models employed.
-Lower precision LLMs can sometimes outperform their higher-precision counterparts, and quantized models may even surpass full-precision models, particularly in the case of larger LLMs.
-Overall, integer quantization (bib8) yields higher-quality explanations compared to other quantization techniques across the selected explainability and interpretability methods.
-Furthermore, a user study involving CFEs and NLEs reveals that explanations generated by quantized LLMs tend to be less coherent and trustworthy.
-Additionally, automatic evaluation metrics for selected explainability methods are not entirely reliable, frequently overestimating the impact of quantization on explanation quality.
-Limitations
-Our experimental work is confined to English-language datasets.
-Consequently, the effectiveness in other languages may not be comparable.
-Extending experiments to the multilingual setting is considered as future work.
-In our experiments, we extensively compare full-precision models with different quantized versions in 4-bit and 8-bit formats.
-Lower-bit quantization, such as 1-bit or 2-bit, is not included in our study.
-Moreover, the scope of our experiments is limited to post-training quantization techniques.
-Investigating the impact of weight-activation quantization, KV cache compression, or quantization-aware training techniques on explainability and interpretability is counted as future work.
-As our experiments are limited to two explainability and two interpretability methods (§3.1), the findings may not fully generalize to all other explainability and interpretability approaches.
-We consider our work to be a first step at the emerging intersection between explainability and interpretability with model efficiency, and extending this analysis to a broader range of methods constitutes a valuable direction for future research within the community.
-Ethics Statement
-The participants in our user studies were compensated at or above the minimum wage in accordance with the standards of our host institutions’ regions.
-The annotation took each annotator 45 minutes on average.
-Appendix A Dataset Information
-A.1 Dataset Examples
-eSNLI
-Figure 6 presents an example from the eSNLI dataset consisting of premise, hypothesis, and gold label.
-LRE
-An exemplary query consists of $s\leftarrow\text{"{the company that created Visual Basic}"}$ and $o\leftarrow\text{"{The current CEO of}"}$.
-A correct answer (in this example: Satya Nadella) by the explained model is the criterion by which the LRE data is filtered.
-TwoHop-Fact
-Figure 7 illustrates an example from the TwoHop-Fact dataset.
-Based on the input text, $(e_{1},r_{1},e_{2})$ corresponds (Superstition, singer, Stevie Wonder), while $(e_{2},r_{2},e_{3})$ represents (Stevie Wonder, mother, Lula).
-eSNLI (Natural Language Inference) Premise: This church choir sings to the masses as they sing joyous songs from the book at a church.
-Hypothesis: The church has cracks in the ceiling.
-Label: Neutral
-Figure 6: An example from the eSNLI dataset.
-TwoHop-Fact (Multi-hop Reasoning) One-hop prompts: ($p_{1}$): The mother of Stevie Wonder is Lula. ($p_{2}$): The singer of ‘Superstition’ is Stevie Wonder.
-Two-hop prompt: ($p$): The mother of the singer of ‘Superstition’ is
-Answer: Lura
-Figure 7: An example from the TwoHop-Fact dataset.
-A.2 Label Distribution
-Label distribution of eSNLI is shown in Figure 8.
-Figure 8: Label distribution of eSNLI.
-Appendix B Models & Inference Time
-Name Citation Size Precision Link
-Llama3 AI@Meta (2024) 8B full https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct
-Llama3 AI@Meta (2024) 8B GPTQ4 https://huggingface.co/TechxGenus/Meta-Llama-3-8B-Instruct-GPTQ
-Llama3 AI@Meta (2024) 8B AWQ https://huggingface.co/TechxGenus/Meta-Llama-3-8B-Instruct-AWQ
-Llama3 AI@Meta (2024) 70B full https://huggingface.co/meta-llama/Meta-Llama-3-70B-Instruct
-Llama3 AI@Meta (2024) 70B GPTQ4 https://huggingface.co/TechxGenus/Meta-Llama-3-70B-Instruct-GPTQ
-Llama3 AI@Meta (2024) 70B AWQ https://huggingface.co/TechxGenus/Meta-Llama-3-70B-Instruct-AWQ
-Qwen2.5 Qwen et al. (2024) 7B Full https://huggingface.co/Qwen/Qwen2.5-7B-Instruct
-Qwen2.5 Qwen et al. (2024) 7B AWQ https://huggingface.co/Qwen/Qwen2.5-7B-Instruct-AWQ
-Qwen2.5 Qwen et al. (2024) 7B GPTQ4 https://huggingface.co/Qwen/Qwen2.5-7B-Instruct-GPTQ-Int4
-Qwen2.5 Qwen et al. (2024) 7B GPTQ8 https://huggingface.co/Qwen/Qwen2.5-7B-Instruct-GPTQ-Int8
-Qwen2.5 Qwen et al. (2024) 14B Full https://huggingface.co/Qwen/Qwen2.5-14B-Instruct
-Qwen2.5 Qwen et al. (2024) 14B AWQ https://huggingface.co/Qwen/Qwen2.5-14B-Instruct-AWQ
-Qwen2.5 Qwen et al. (2024) 14B GPTQ4 https://huggingface.co/Qwen/Qwen2.5-14B-Instruct-GPTQ-Int4
-Qwen2.5 Qwen et al. (2024) 14B GPTQ8 https://huggingface.co/Qwen/Qwen2.5-14B-Instruct-GPTQ-Int8
-Qwen2.5 Qwen et al. (2024) 32B Full https://huggingface.co/Qwen/Qwen2.5-32B-Instruct
-Qwen2.5 Qwen et al. (2024) 32B AWQ https://huggingface.co/Qwen/Qwen2.5-32B-Instruct-AWQ
-Qwen2.5 Qwen et al. (2024) 32B GPTQ4 https://huggingface.co/Qwen/Qwen2.5-32B-Instruct-GPTQ-Int4
-Qwen2.5 Qwen et al. (2024) 32B GPTQ8 https://huggingface.co/Qwen/Qwen2.5-32B-Instruct-GPTQ-Int8
-Qwen2.5 Qwen et al. (2024) 72B Full https://huggingface.co/Qwen/Qwen2.5-72B-Instruct
-Qwen2.5 Qwen et al. (2024) 72B AWQ https://huggingface.co/Qwen/Qwen2.5-72B-Instruct-AWQ
-Qwen2.5 Qwen et al. (2024) 72B GPTQ4 https://huggingface.co/Qwen/Qwen2.5-72B-Instruct-GPTQ-Int4
-Qwen2.5 Qwen et al. (2024) 72B GPTQ8 https://huggingface.co/Qwen/Qwen2.5-72B-Instruct-GPTQ-Int8
-Table 5: Detailed information about used LLMs in our experiments.
-Table 5 presents details of the all LLMs used in our experiments (§3.4), including model sizes, quantization approaches and corresponding URLs from the Hugging Face Hub.
-All models were directly obtained from the Hugging Face repository.
-All experiments were conducted using A100 or H100 GPUs.
-Explanation generation across the entire dataset, including both natural language explanations (NLEs) and counterfactual examples (CFEs), can be completed within 10 hours.
-Appendix C Additional Experiments
-C.1 Models
-In addition to the LLMs listed in Section 3.4, we further validate our findings using models from a different family, specifically Llama3-8B and Llama3-70B AI@Meta (2024).
-Detailed information about these two models is provided in Table 5.
-C.2 Datasets
-Additional experiments further evaluate counterfactual examples performance on AG News Zhang et al. (2015) and natural language explanations on HealthFC Vladika et al. (2024).
-C.2.1 Dataset Information
-AG News
-Zhang et al. (2015) comprises news articles generated by merging the title and description fields from articles across four categories: World, Sports, Business, and Sci/Tech.
-HealthFC
-Vladika et al.
-(2024) is a bilingual fact-checking dataset (English and German) comprising questions, documents, veracity annotations (indicating whether the answer is true, false, or unknown based on the provided document), and corresponding human-annotated explanations.
-C.2.2 Dataset Examples
-Examples from the AG News and HealthFC datasets are shown in Figure 9.
-AG News (News Topic Classification) News: E-mail scam targets police chief Wiltshire Police warns about ""phishing"" after its fraud squad chief was targeted.
-Label: sci/tech
-HealthFC (Fact-checking) Question: Does chicken soup help with colds?
-Document: "However, such experiments – even though they may sound so promising – do not provide any evidence that the soup also works the same in the human body.
-Chicken soup in case of cold: theories without evidence Extensive studies on the effect of chicken soup do not exist, but all the more attempts to explain them.
-As long as there are no studies with human subjects, we simply cannot assess whether and which chicken soup ingredients could help cold-stricken people."
-Label: unknown
-Explanation: "So far, this has only been investigated in laboratory experiments.
-Studies on efficacy in humans are missing so far.
-Therefore, we can not judge whether chicken soup is helpful for colds."
-Figure 9: Examples from AG News and HealthFC.
-C.2.3 Label Distributions
-Figure 10 demonstrates the label distributions of the AG News and HealthFC datasets.
-(a) AG News
-(b) HealthFC
-Figure 10: Label distributions of AG News and HealthFC.
-C.3 Natural Language Explanation
-Model Dataset eSNLI HealthFC
-Precision $F_{1}$ BART TIGER $F_{1}$ BART TIGER
-Score $\uparrow$ Score $\uparrow$ Score $\uparrow$ Score $\uparrow$ Score $\uparrow$ Score $\uparrow$
-Llama3-8B full 36.90% -6.62 -0.37 51.43% -4.29 -1.01
-bib4 34.40% -6.60 -0.54 54.00% -4.65 -1.30
-bib8 37.60% -6.66 -0.37 51.14% -4.26 -0.95
-gptq4 38.70% -7.42 -0.29 57.71% -4.36 -1.28
-awq 35.70% -6.76 -0.26 56.00% -4.43 -1.31
-Llama3-70B full 57.80% -6.58 -0.16 44.00% -4.48 -1.04
-bib4 59.30% -6.62 -0.71 55.71% -4.44 -1.56
-bib8 43.90% -6.76 -0.26 42.86% -4.18 -1.30
-gptq4 60.90% -6.62 -0.14 45.43% -4.32 -1.12
-awq 61.10% -6.61 -0.15 47.71% -4.54 -1.01
-Table 6: Automatic evaluation results of NLEs generated using ZeroCoT by Llama3-8B and Llama3-70B with full precision, AWQ, bib, and GPTQ quantization on eSNLI and HealthFC, assessed using Micro $F_{1}$ score, BARTScore, and TIGERScore.
-Model Dataset HealthFC
-Precision $F_{1}$ BART TIGER
-Score $\uparrow$ Score $\uparrow$ Score $\uparrow$
-Qwen2.5-7B full 39.14% -4.41 -0.34
-bib4 36.57% -4.35 -0.64
-bib8 36.57% -7.84 -0.94
-gptq4 38.86% -4.29 -0.65
-gptq8 39.71% -6.02 -0.74
-awq 36.86% -4.26 -0.93
-Qwen2.5-14B full 45.14% -4.25 -0.39
-bib4 40.29% -4.30 -0.29
-bib8 40.29% -4.34 -0.24
-gptq4 36.57% -4.23 -0.41
-gptq8 39.42% -4.27 -0.31
-awq 46.29% -4.33 -0.48
-Qwen2.5-32B full 42.57% -8.17 -2.68
-bib4 42.86% -6.06 -1.61
-bib8 47.43% -9.07 -3.00
-gptq4 41.43% -6.62 -1.88
-gptq8 42.29% -8.54 -2.86
-awq 39.43% -8.97 -3.27
-Qwen2.5-72B full 51.43% -4.21 -0.58
-bib4 53.71% -4.21 -0.58
-bib8 54.00% -4.22 -0.57
-gptq4 53.42% -4.17 -0.60
-gptq8 52.09% -4.20 -0.65
-awq 52.28% -4.22 -0.66
-Table 7: Automatic evaluation results of natural language explanations (NLEs) generated using ZeroCoT by Qwen2.5-{7,14,32,72}B with full precision and AWQ, bib, and GPTQ $\{4,8\}$-bits quantization on HealthFC, assessed by Micro $F_{1}$ score, BARTScore, and TIGERScore.
-Table 6 presents automatic evaluation results of natural language explanations generated by Llama3-{8B,70B} on the eSNLI and HealthFC datasets.
-Table 7 provides additional evaluation results of natural language explanations generated by Vicuna-13B, and Qwen2.5-{7B,14B,32B,72B} on the HealthFC dataset.
-Notably, under some circumstances, quantized models outperform their full-precision counterparts.
-Furthermore, both tables indicate that in general, larger quantized models exhibit superior performance compared to smaller full-precision models.
-Additionally, integer quantization implemented by BitsAndBytes and GPTQ demonstrate compelling performance on average compared to AWQ.
-C.4 Counterfactual Example
-Model Dataset eSNLI AG News
-Precision LFR $\uparrow$ PPL $\downarrow$ TS $\downarrow$ LFR $\uparrow$ PPL $\downarrow$ TS $\downarrow$
-Llama3-8B full 90.40% 62.26 1.83 48.80% 42.88 1.54
-bib4 82.00% 81.80 1.01 48.40% 46.01 1.38
-bib8 86.80% 72.17 1.80 51.60% 46.95 1.60
-gptq4 88.20% 76.55 1.99 65.20% 62.61 1.60
-awq 87.40% 75.21 1.93 54.00% 79.87 1.21
-Llama3-70B full 71.40% 80.77 0.75 49.00% 101.97 0.40
-bib4 88.40% 139.80 0.58 49.60% 130.05 0.48
-bib8 75.93% 102.23 1.70 26.20% 140.24 0.25
-gptq4 90.40% 94.48 0.69 53.20% 115.03 0.52
-awq 91.20% 114.89 0.51 47.40% 106.90 0.42
-Table 8: Automatic evaluation results of CFEs generated using FIZLE by Llama3-8B and Llama3-70B with full precision, AWQ, bib, and GPTQ quantization on eSNLI and HealthFC, assessed using label flip rate (LFR), perplexity (PPL), and text similarity (TS).
-Model Dataset AG News
-Precision LFR $\uparrow$ PPL $\downarrow$ TS $\downarrow$
-Qwen2.5-7B full 35.40% 74.87 0.53
-bib4 36.00% 95.26 0.61
-bib8 42.00% 72.97 0.61
-gptq4 34.80% 81.52 0.57
-gptq8 31.20% 79.45 0.57
-awq 38.40% 80.52 0.57
-Qwen2.5-14B full 38.40% 102.83 0.66
-bib4 42.20% 133.47 0.78
-bib8 34.60% 99.28 0.63
-gptq4 39.40% 99.97 0.66
-gptq8 36.80% 95.11 0.66
-awq 39.20% 84.55 0.61
-Qwen2.5-32B full 31.40% 83.13 0.49
-bib4 39.20% 80.93 0.56
-bib8 34.00% 82.82 0.49
-gptq4 33.40% 81.48 0.48
-gptq8 35.40% 84.81 0.49
-awq 37.20% 80.60 0.52
-Qwen2.5-72B full 28.80% 96.45 0.43
-bib4 25.40% 93.03 0.45
-bib8 31.20% 91.51 0.44
-gptq4 29.60% 96.05 0.44
-gptq8 26.40% 96.47 0.45
-awq 24.60% 88.97 0.42
-Table 9: Automatic evaluation results of CFEs generated using FIZLE by Qwen2.5-{7,14,32,72}B with full precision and AWQ, bib, and GPTQ quantization on HealthFC, assessed by label flip rate (LFR), perplexity (PPL), and text similarity (TS).
-Table 8 displays the automatic evaluation results for counterfactuals generated by Llama3-{8B,70B}, assessed on the eSNLI and AG News datasets.
-Table 9 supplements this by presenting further evaluation results for counterfactual examples produced by Qwen2.5-{7B,14B,32B,72B}, evaluated on the AG News dataset.
-Table 8 and Table 9 indicate that the quality of counterfactual examples does not uniformly enhance with larger model sizes, a finding consistent with observations shown by Nguyen et al.
-(2024); Bhattacharjee et al.
-(2024); Li et al.
-(2024d).
-We find that occasionally full-precision models outperform quantized models in generating effective counterfactual examples on AG News and eSNLI.
-Generally, integer quantization implemented by BitsAndBytes demonstrates superior performance compared to AWQ and GPTQ.
-Moreover, the conclusions of Badshah and Sajjad (2024) are not corroborated; as evidenced in Table 8 and Table 9, larger quantized models sometimes exhibit inferior performance compared to smaller full-precision models.
-C.5 Knowledge Memorization Analysis
-Knowledge Recall Accuracy.
+Table 3 shows the latent multi-hop reasoning accuracy comparison between full-precision models and quantized models.
+Additionally, Figure 12, Figure 14, and Figure 15 display the differences in the entity recall score, consistency score, and accuracy between the AWQ, GPTQ8, GPTQ4, bib8, bib4 quantized and full-precision models of Qwen2.5-7B, Qwen2.5-14B, and Llama3-8B, evaluated across all layers.
+(a) AWQ
+(b) GPTQ8
+(c) GPTQ4
+(d) bib8
+(e) bib4
+Figure 12: Difference in the entity recall score (EntRec), consistency score (CnstScore), and accuracy between the AWQ, GPTQ8, GPTQ4, bib8, bib4 quantized and full-precision models of Qwen2.5-7B, evaluated across all layers.
+(a) AWQ
+(b) GPTQ8
+(c) GPTQ4
+(d) bib8
+(e) bib4
+Figure 14: Difference in the entity recall score (EntRec), consistency score (CnstScore), and accuracy between the AWQ, GPTQ8, GPTQ4, bib8, bib4 quantized and full-precision models of Qwen2.5-14B, evaluated across all layers.
+(a) AWQ
+(b) GPTQ
+(c) bib8
+(d) bib4
+Figure 15: Difference in the entity recall score (EntRec), consistency score (CnstScore), and accuracy between the AWQ, GPTQ, bib8, bib4 quantized and full-precision models of Llama3-8B, evaluated across all layers.
 Qwen2.5-7B full bib4 bib8 gptq4 gptq8 awq
@@
 Table 5: Per-relation knowledge recall accuracy results (%) on the LRE dataset for Qwen2.5-14B across different quantization methods.
-Qwen2.5-32B full bib4 bib8 gptq4 gptq8 awq
-city in country 85.00 85.00 85.00 85.00 85.00 85.00
-company CEO 57.00 55.00 55.00 57.00 57.00 57.00
-company hq 68.00 67.00 68.00 66.00 67.00 67.00
-country capital city 100.00 100.00 100.00 100.00 100.00 100.00
-country currency 97.00 97.00 93.00 93.00 93.00 93.00
-country language 96.00 96.00 96.00 96.00 96.00 96.00
-country largest city 100.00 100.00 100.00 100.00 100.00 100.00
-food from country 77.00 77.00 77.00 77.00 77.00 77.00
-landmark in country 87.00 86.00 86.00 87.00 87.00 87.00
-landmark on continent 86.00 84.00 86.00 86.00 86.00 82.00
-person father 65.00 61.00 65.00 60.00 65.00 63.00
-person lead singer of band 100.00 95.00 100.00 95.00 100.00 100.00
-person mother 59.00 56.00 59.00 54.00 59.00 57.00
-person native language 94.00 93.00 94.00 94.00 94.00 94.00
-person occupation 49.00 47.00 51.00 48.00 50.00 49.00
-person plays instrument 62.00 61.00 62.00 62.00 62.00 63.00
-person sport position 68.00 64.00 67.00 68.00 67.00 68.00
-person university 48.00 48.00 51.00 46.00 49.00 46.00
-plays pro sport 88.00 85.00 89.00 88.00 88.00 85.00
-pokemon evolution 100.00 100.00 100.00 100.00 100.00 100.00
-president birth year 0.00 0.00 0.00 0.00 0.00 0.00
-president election year 0.00 0.00 0.00 0.00 0.00 0.00
-product by company 82.00 82.00 82.00 83.00 82.00 81.00
-star constellation name 89.00 87.00 89.00 88.00 89.00 89.00
-superhero archnemesis 45.00 43.00 46.00 42.00 45.00 43.00
-superhero person 72.00 69.00 70.00 76.00 72.00 74.00
-AVG 72.45 70.47 72.40 71.25 72.39 71.47
-Table 12: Per-relation knowledge recall accuracy results (%) on the LRE dataset for Qwen2.5-32B across different quantization methods.
-Qwen2.5-72B full bib4 bib8 gptq4 gptq8 awq
-city in country 96.00 100.00 96.00 96.00 96.00 96.00
-company CEO 62.00 62.00 63.00 61.00 62.00 61.00
-company hq 73.00 73.00 73.00 73.00 73.00 73.00
-country capital city 100.00 100.00 100.00 100.00 100.00 100.00
-country currency 80.00 80.00 80.00 80.00 80.00 83.00
-country language 96.00 96.00 96.00 96.00 96.00 96.00
-country largest city 100.00 100.00 100.00 100.00 100.00 100.00
-food from country 83.00 83.00 83.00 83.00 83.00 83.00
-landmark in country 89.00 89.00 90.00 89.00 89.00 89.00
-landmark on continent 93.00 92.00 93.00 92.00 93.00 93.00
-person father 74.00 71.00 74.00 69.00 74.00 71.00
-person lead singer of band 90.00 100.00 90.00 95.00 90.00 100.00
-person mother 69.00 65.00 69.00 64.00 70.00 66.00
-person native language 93.00 93.00 93.00 92.00 93.00 93.00
-person occupation 56.00 54.00 55.00 56.00 56.00 53.00
-person plays instrument 73.00 69.00 73.00 71.00 73.00 69.00
-person sport position 78.00 75.00 78.00 76.00 78.00 73.00
-person university 52.00 49.00 53.00 49.00 52.00 52.00
-plays pro sport 93.00 93.00 93.00 93.00 93.00 93.00
-pokemon evolution 100.00 100.00 100.00 100.00 100.00 100.00
-president birth year 0.00 0.00 0.00 0.00 0.00 0.00
-president election year 0.00 0.00 0.00 0.00 0.00 0.00
-product by company 85.00 85.00 85.00 85.00 85.00 84.00
-star constellation name 88.00 89.00 89.00 88.00 88.00 88.00
-superhero archnemesis 48.00 43.00 49.00 48.00 48.00 49.00
-superhero person 78.00 80.00 78.00 79.00 78.00 77.00
-AVG 78.26 76.76 78.22 76.77 78.32 76.72
-Table 13: Per-relation knowledge recall accuracy results (%) on the LRE dataset for Qwen2.5-72B across different quantization methods.
-In Table 10, 11, 12, 13, we provide the per-relation factual recall accuracy on Qwen2.5-{7B, 14B, 32B, 72B} models, respectively.
-As discussion in Section 5.1.2, we observe that the accuracy drop is more severe in relations where the original performance has not yet saturated.
-Neuron Attribution Analysis.
-Here we present the neuron-level knowledge attribution results on relations person father, person mother, person sport position, complementing our results in Section 5.1.2.
-Analysis on layer-wise neuron contribution scores are shown in Figure 11, the analysis on top-300 neurons are given in Figure 12.
-(c) Attention: Person father
-(d) FFN: Person father
-(e) Attention: Person mother
-(f) FFN: Person mother
-(g) Attention: Person sport position
-(h) FFN: Person sport position
-Figure 11: Contribution score drop across quantization methods for other relationship types, comparing attention sublayers (left) and feed-forward sublayers (right).
-Rows show different relationships: (a-b) person father, (c-d) person mother, and (e-f) person sport position.
-The plots reveal how different quantization methods affect knowledge representation across model layers.
-Refer to caption Refer to caption
-(a) Person father: Contribution scores (b) Person father: Layer distribution
-Refer to caption Refer to caption
-(c) Person mother: Contribution scores (d) Person mother: Layer distribution
-Refer to caption Refer to caption
-(e) Person sport position: Contribution scores (f) Person sport position: Layer distribution
-Figure 12: Analysis of other relationship types under different quantization methods.
-Left column: Average contribution scores of top 300 feed-forward neurons across quantization methods.
-Right column: Distribution of high-scoring neurons across model layers.
-Each row represents a different relationship: (a-b) person father, (c-d) person mother, and (e-f) person sport position.
-These visualizations reveal both the magnitude of contribution score changes and their distribution across the model architecture when applying different quantization techniques to various types of factual knowledge.
-Appendix D Annotation
-Figure 13 displays annotation guideline that we provide to human annotators.
-CFEs and NLEs are presented to annotators in the form of questionnaires.
-We use the Crowdee [^7] https://www.crowdee.com/ crowdsourcing platform to recruit annotators, distribute the questionnaires, and store their responses.
-A total of 48 annotators were recruited, all of whom are native English speakers without requiring specific expertise in explainable AI (XAI).
-Each annotators will be given 15 explanations, along with two evaluation dimensions (§4.2.1).
-Each explanation will be evaluated by at least two annotators.
-Annotation Guideline for Human Evaluation ### User Study Description: Dear participants, Thanks for attending our user study.
-This study focuses on evaluating model-generated explanations.
-We present two types of explanations: Counterfactual Example (CFE): A minimally edited version of the input text that results in a change in the model’s prediction.
-Natural Language Explanation (NLE): A textual justification generated by the model to explain its decision-making process for a given input.
-Each explanation should be evaluated along the following two dimensions.
-Please assign a score to each dimension on a scale from 1 (strongly disagree) to 5 (strongly agree).
-Trustworthiness: Evaluate whether the provided explanation is truthfulness and can be relied upon by humans; Coherence: Assess whether the provided explanation is sensible, clear, and coherent, and effectively captures the rationale;
-### Dataset Structure: e-SNLI (Stanford Natural Language Inference): Each example consists of a premise and a hypothesis.
-The task is to determine the relationship between the two, categorizing it as either Entailment, Contradiction, or Neutral based on the information in the premise.
-e-SNLI Example: {example}
-Explanation: {example explanation}
-Rating:
-Trustworthiness: {score}
-Coherence: {score}
-Entailment means the hypothesis must be true if the premise is true.
-Contradiction means the hypothesis must be false if the premise is true.
-Neutral means the hypothesis might be true, or might not — we can’t tell just from the premise.
-### User Study Instruction: You will be provided with 15 instances to evaluate.
-For the counterfactual example evaluation, each instance includes a premise–hypothesis pair.
-Your task is to evaluate only the quality of the premise according to the two dimensions described above.
-For the natural language explanation evaluation, each instance also includes a premise–hypothesis pair, along with a model-generated justification in natural language.
-In this case, your task is to evaluate the provided justification based on the same two dimensions.
-Figure 13: Annotation Guideline.
-Appendix E Performance Degradation Comparison
-Figure 14 illustrates the performance degradation between automatic and human evaluations, averaged over three automatic metrics and three human evaluation dimensions.
-(a) Counterfactual Example Evaluation on Qwen2.5-7B.
-(b) Counterfactual Example Evaluation on Qwen2.5-32B.
-(c) Counterfactual Example Evaluation on Qwen2.5-72B.
-(d) Natural Language Explanation Evaluation on Qwen2.5-7B.
-(e) Natural Language Explanation Evaluation on Qwen2.5-32B.
-(f) Natural Language Explanation Evaluation on Qwen2.5-72B.
-Figure 14: Performance degradation between automatic and human evaluations, averaged across three automatic evaluation metrics and three human evaluation dimensions.